### v2

# Safe MPC Alignment with Human Directional Feedback

Abstract
In safety-critical robot planning or control, manually specifying safety constraints or learning them from demonstrations can be challenging. In this article, we propose a certifiable alignment method for a robot to learn a safety constraint in its model predictive control (MPC) policy with human online directional feedback. To our knowledge, it is the first method to learn safety constraints from human feedback. The proposed method is based on an empirical observation: human directional feedback, when available, tends to guide the robot toward safer regions. The method only requires the direction of human feedback to update the learning hypothesis space. It is certifiable, providing an upper bound on the total number of human feedback in the case of successful learning, or declaring the hypothesis misspecification, i.e., the true implicit safety constraint cannot be found within the specified hypothesis space. We evaluated the proposed method using numerical examples and user studies in two simulation games. Additionally, we implemented and tested the proposed method on a real-world Franka robot arm performing mobile water-pouring tasks. The results demonstrate the efficacy and efficiency of our method, showing that it enables a robot to successfully learn safety constraints with a small handful (tens) of human directional corrections.
Index Terms:
Learning from Human Feedback, Constraint Inference, Model Predictive Control, Human-Robot Interaction, Cutting-Plane Method.
I Introduction
In robot policy design, safety is always the top priority, especially in safety-critical applications, such as autonomous vehicles or robots operating near humans, where safety violation can lead to irreversible loss or damage. While extensive work in safe control [1] and reinforcement learning [2] focuses on designing safe policy given robot motion constraints, less attention has been paid to how to specify the safety constraints in the first place. In many applications, specifying proper safety constraints can be challenging. First, explicit safety boundary for robot behavior can be hard to model. For example, when delivering food in a crowded restaurant, a service robot must balance the tray to avoid spills; this requires constraints on robot speed, food weight distributions, and interaction with its surroundings, and all those constraint aspects can be complex to model. Second, safety perception can vary among users. In the same example of the delivery robot: some diner may feel secure with the robot staying within a foot or two when delivering items, while others may feel uncomfortable with close proximity and prefer the robot to stay at a slight distance.
Figure 1: Overall illustration of the proposed safe MPC alignment. Our method online updates the safety constraint in robot safe MPC policy based on human directional feedback, such that the robot eventually aligns the safe MPC policy with human intent for accomplishing safety-critical tasks.
The challenges in safety specifications lead to two questions, which this paper seeks to answer. Q1: can a robot learn and align its safety constraints with human intent in an interactive manner? Q2: can such learning process to be user-data efficient such that a robot can quickly adapt to different users? Recent work [3, 4, 5] has explored learning safety constraints from human demonstrations. However, acquiring safety-informative demonstrations could be challenging. An alternative is learning from human preference feedback [6, 7, 8, 9, 10, 11, 12], where users express preferences between trajectories to guide robot learning. Despite the success in reward alignment [6], the query trajectories can be difficult for humans to judge when they are equally suboptimal, especially in early learning stages. To address such issue, [13, 14, 15, 16, 17, 18, 19] proposed learning from human physical corrections. While promising, the recent work [20] suggests those methods are potentially sensitive to correction magnitude: improper human corrections can lead to overshooting and inefficiency of learning. Importantly, all those feedback learning methods primarily focus on learning the reward function and do not address the challenges of learning safety constraints from human feedback.
This paper aims to answer the previous two questions by proposing a new paradigm for a robot to efficiently learn safety constraints in its model predictive control (MPC) policy with human feedback. We name the method safe MPC alignment. As shown in Fig. 1, a human user, observing the robot’s potentially unsafe behavior, provides feedback to the robot. Upon receiving human feedback, the robot immediately updates the safety constraints in its MPC policy and continues with the updated policy. Different from existing learning from feedback methods, our method only requires human directional corrections [20], which can be sparse in time. For example, when a human user corrects an autonomous car, he/she only needs to give the command of ‘turn left‘ or ‘turn right‘, without dictating how much the car needs to turn. The safe MPC alignment only utilizes the direction of human correction, regardless of the magnitude, for the safety constraints update. As shown in our previous work [20] or later, the type of directional feedback provides a consistently large and flexible correction space, effectively overcoming the over-shooting correction, compared to existing magnitude feedback. Importantly, the proposed alignment is able to guarantee the exponential convergence with finite rounds of human corrections, or output failure with certified hypothesis misspecification. Specifically, we highlight our main contributions below.
(I) We propose the method of safe MPC alignment with human directional feedback. The method enables a robot to online learn a safe MPC policy from human directional corrections. To our knowledge, it is the first work for interactive learning of safety constraints from human feedback.
(II) We theoretically show the human-data efficiency of the safe MPC alignment: the proposed method is guaranteed to have exponential convergence with finite human corrections. Furthermore, we prove that the proposed method enables the certification of hypothesis space misspecification failure.
(III) We conduct extensive user studies to validate the proposed method, including two MuJoCo-based simulation games and a real-world experiment of teaching a Franka robot arm for mobile liquid pouring. The results show the effectiveness and efficiency of our method.
The remaining paper is organized as follows. Section II reviews the related work, and Section III formulates the safe MPC alignment problem. The main method is presented in Section IV. Section V gives the implementation details and presents the main results. Sections VI and VII provide the experiments and user studies. Conclusions are in Section VIII.
II Related Work
II-A Learning from Demonstrations
Under the general theme of learning from human data, the proposed method is loosely connected to the topic of learning from demonstrations, where cost/reward functions are learned from expert (human) data using inverse reinforcement learning (IRL) and inverse optimal control (IOC) methods [21, 22]. IRL/IOC methods typically run offline: human demonstration data is first obtained, and then reward/cost learning is performed [22, 23, 24, 25, 26, 27]. As pointed in [28], acquiring optimal demonstration data can be difficult in practice. The challenge can be either attributed to the higher degrees of freedom of the robot, or the lack of prior knowledge about the desired robot behaviour to perform demonstrations with high-quality [29].
In this work, we focus on interactive learning of safety constraints. Different from learning from demonstrations that needs the exhibition of desired robot motion, the proposed method only requires a human user to incrementally improve the current robot motion by providing directional feedback, thus more accessible to human users.
II-B Learning from Human Feedback
The line of this work typically focuses on learning rewards/costs from human feedback, which can be given in preference and corrections. In learning from human preferences, human preference is typically given as user’s ranking of two or multiple robot candidate trajectories [6, 7, 8, 9, 10, 11, 12], and Bradley-Terry model [30] is used to update the rewards/costs. The recent variant seeks to directly learn a policy from human preference [31].
Human feedback can also come as corrections, where human users may intervene and correct the robot at runtime. An intended robot reward/cost function can be inferred from human corrections. Specific human corrections include language [32, 33, 34, 35, 36] and physical corrections [13, 15, 18]. Since human corrections can be sparse in time, many methods [13, 14, 15, 16, 17, 18, 19] use trajectory deformation methods to interpolate sparse corrections as a continuous trajectory. The rewards/costs are updated to explain the deformed trajectories. Some strategies are further proposed to deal with the uncertainty and improve the robustness of the learning process, like updating one parameter dimension at a time [15] or formulating the reward/cost learning in the framework of Kalman Filter [14].
Our work belongs to the category of learning from human corrections. But instead of learning rewards/costs, we focus on learning safety constraints from human corrections, which has not been addressed before. Methodologically, we propose a certifiable method, which guarantees the exponential learning convergence with finite human corrections, and certifies the hypothesis misspecification if the learning hypothesis space does not contain the true human intent. Those theoretical results are missing in existing methods.
II-C Learning Constraints from Demonstrations
An increasing number of work focuses on learning safety constraints from demonstrations. One vanilla idea is to formulate constraint learning as a binary classification. However, since a robot usually runs within a safe region, negative demonstration samples that violate safety are challenging to collect. To address such, [3, 37] consider the trajectories of lower cost as the unsafe samples. [38] proposes adjusting the constraints toward decreasing the safety margin of feasible policies. Safety constraints can also be learned via mixed integer programming [3, 37], pessimistic adjustment [38], and no-regret algorithm [39].
Another line of work attempts to learn safety constraints only from safe demonstration data. For example, some work exploits the optimality of demonstration with unknown constraints [40, 41, 42]. Different probabilistic formulations are also proposed to learn constraints from demonstration data, such as maximum likelihood [5, 43, 44], maximum causal entropy [45, 46], and Bayesian inference [47, 48]. In the above constraint learning, safety constraints can be represented as a set of safe states [40], the belief of potential constraints [41], Gaussian process [42], or directly as a convex hull in policy space [4].
The vast majority of the above methods formulate the constraint learning problem in the setting of reinforcement learning, and thus the learning process usually involves repetitive policy optimization. Those methods cannot be deployed in an online setting, where a robot is expected to on-the-fly update its safe policy and respond to human feedback. To fill this gap, we will develop the first method, to our best knowledge, to enable online learning of safe constraint from human feedback.
III Preliminary and Problem Formulation
III-A Safe Model Predictive Control
We consider the following safe model predictive control (Safe MPC) for a robotic system, where the safety constraint is learnable and parameterized by $\boldsymbol{\theta}\in\mathbb{R}^{r}$.
$\min_{\boldsymbol{u}_{0:T-1}}&\quad J(\boldsymbol{u}_{0:T-1};\boldsymbol{x}_{0})=\sum_{t=0}^{T-1}c(\boldsymbol{x}_{t},\boldsymbol{u}_{t})+h(\boldsymbol{x}_{T})\\ \text{subject to}&\quad\boldsymbol{x}_{t+1}=\boldsymbol{f}(\boldsymbol{x}_{t},\boldsymbol{u}_{t})\\ &\quad{g}_{\boldsymbol{\theta}}(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1})\leq 0\end{aligned}.$ (1)
Here, $\boldsymbol{x}_{t}\in\mathbb{R}^{n}$ and $\boldsymbol{u}_{t}\in\mathbb{R}^{m}$ are robot state and action, respectively; $t=0,1,2,...,T$ is the time step with prediction horizon $T$; $\boldsymbol{x}_{0:T}=(\boldsymbol{x}_{0},\boldsymbol{x}_{1},...,\boldsymbol{x}_{T})$ and $\boldsymbol{u}_{0:T-1}=(\boldsymbol{u}_{0},\boldsymbol{u}_{1},...,\boldsymbol{u}_{T-1})$; $J(\cdot)$ is the given cost function including the running cost ${c}(\cdot)$ and final cost $h(\cdot)$, representing the given task; $\boldsymbol{f}(\cdot,\cdot)$ is the system dynamics; ${g}_{\boldsymbol{\theta}}(\cdot)$ is the safety constraint, which is parameterized by $\boldsymbol{\theta}\in\mathbb{R}^{r}$. Given starting state $\boldsymbol{x}_{0}$ and constraint parameter $\boldsymbol{\theta}$, the trajectory solution to (1), denoted as $\boldsymbol{\xi}_{\boldsymbol{\theta}}=\{\boldsymbol{u}^{\boldsymbol{\theta}}_{0:T-1},\boldsymbol{x}^{\boldsymbol{\theta}}_{0:T}\}$, is an implicit function of $\boldsymbol{\theta}$.
In a receding horizon implementation, $\boldsymbol{x}_{0}$ is set as the robot’s current state (estimation) $\boldsymbol{{x}}_{k}$, where $k=0,1,2,...,$ is the policy rollout time step; and only the first action $\boldsymbol{u}^{\boldsymbol{\theta}}_{0}$ in $\boldsymbol{{u}}_{0:T}^{\boldsymbol{\theta}}$ is applied to the robot to move the robot to the next state $\boldsymbol{x}_{k+1}$. This leads to a closed-loop Safe MPC policy for the robot, denoted as $\boldsymbol{\pi}_{\boldsymbol{\theta}}$, mapping from robot current state $\boldsymbol{x}_{k}$ to its action $\boldsymbol{u}^{\boldsymbol{\theta}}_{0}$.
III-B Safe Asymptotic Approximation of Safe MPC
We present a penalty-based approximation to the Safe MPC in (1). Define trajectory variable $\boldsymbol{\xi}=\{\boldsymbol{u}_{0:T-1},\boldsymbol{x}_{0:T}\}$ and the following penalty objective:
$B(\boldsymbol{\xi},\boldsymbol{\theta})=J(\boldsymbol{\xi})-\gamma\ln(-g_{\boldsymbol{\theta}}(\boldsymbol{\xi})).$ (2)
Here, the above defined penalty $B(\boldsymbol{\xi},\boldsymbol{\theta})$ includes the original control cost in (1) and safety constraint $g_{\boldsymbol{\theta}}(\boldsymbol{\xi})$ via logarithmic barrier transformation, $\gamma>0$ is named the barrier parameter, controlling the weight of safety penalty $-\ln(-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}))$ relative the control cost. With $B(\boldsymbol{\xi},\boldsymbol{\theta})$, one can establish the following penalty-based MPC formulation
$\min_{\boldsymbol{u}_{0:T-1}}\quad B(\boldsymbol{\xi},\boldsymbol{\theta})\qquad\text{subject to}\quad\boldsymbol{x}_{t+1}=\boldsymbol{f}(\boldsymbol{x}_{t},\boldsymbol{u}_{t}).$ (3)
We denote the solution to the above penalty-based MPC (3) as $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}=\{\boldsymbol{u}_{0:T-1}^{\boldsymbol{\theta},\gamma},\boldsymbol{x}_{0:T-1}^{\boldsymbol{\theta},\gamma}\}$. The following result is from our prior work [49], stating that the penalty-based MPC (3) is an asymptotic and safe approximation to the Safe MPC (1).
Lemma 1.
(Theorem 2 in [49]) If the original Safe MPC (1) satisfies certain mild conditions (i.e., local second-order condition and strict complementarity), then for any small $\gamma>0$, the penalty-based MPC (3) has a (local) solution $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, and $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ is safe with respect to the original constraint, i.e., $g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})<0$, and $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ satisifies the first-order condition of (3). Furthermore, $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ is asymptotically approximate $\xi_{\boldsymbol{\theta}}$: $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}\rightarrow\boldsymbol{\xi}_{\boldsymbol{\theta}}$ as $\gamma\rightarrow 0$.
The above lemma asserts that solving the Safe MPC (1) can be replaced by solving penalty-based MPC (3). The solution approximation is controlled by $\gamma$ for arbitrary accuracy, as $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}\rightarrow\boldsymbol{\xi}_{\boldsymbol{\theta}}$ as $\gamma\rightarrow 0$. A good approximation is achieved by smaller $\gamma$. on the contrary, if $\gamma$ is larger, penalty-based MPC (3) will lead to more safety-conservative solution $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$. In the development below, we fix the parameter $\gamma$ in (3). We denote the penalty-based MPC policy in (3) as $\boldsymbol{\pi}^{\gamma}_{\theta}$.
III-C Problem Formulation
III-C1 Representation of Human Intended Safety
Consider a human-robot coexistence task, where a human is supervising the robot motion. The user has his/her intent of robot motion safety. In Safe MPC (1), we hypothesize user-intended safety can be represented by a subset of the trajectory space:
$\begin{split}\mathcal{S}(\boldsymbol{\theta}_{H})=\{(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1})\,\,|\,\,\ {g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1})\leq 0\}.\end{split}$ (4)
Here, $\boldsymbol{\theta}_{H}$ is intent parameter from an implicit set $\boldsymbol{\theta}_{H}\in\Theta_{H}\subset\mathbb{R}^{r}$. $\Theta_{H}$ is implicit to human and unknown to the robot. For the robot policy $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$, if $\boldsymbol{\theta}\in\Theta_{H}$, the user considers the robot motion to be safe; otherwise if $\boldsymbol{\theta}\notin\Theta_{H}$, the robot motion by $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$ might not be safe at certain time in the view of the user, and thus the user provides corrections to improve its safety. We use an implicit set $\Theta_{H}\in\mathbb{R}^{r}$, instead of single point $\boldsymbol{\theta}_{H}$, to capture the potential variation of user intents. We suppose that a volume measure of $\Theta_{H}$, $\operatorname{{Vol}}(\Theta_{H})$, is non-zero, and a user intent $\boldsymbol{\theta}_{H}\in\Theta_{H}$ can vary at different steps and context, but always within $\Theta_{H}$.
III-C2 Human Directional Correction
We consider human gives correction to a robot in an online fashion while the robot is executing its most updated $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$. Specifically, let $i$ be the index of learning update, and suppose robot current policy is $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$. During rolling out $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, the robot receives a sparse human directional correction $\boldsymbol{a}_{k_{i}}\in\mathbb{R}^{m}$ at the rollout time step $k_{i}$.
Here, we do not assume precise timing $k_{i}$ or duration of the human correction due to the sparsity or intermittency of human feedback. In other words, human can apply one or multiple corrections at any time in robot rollout time steps for any duration as he/she finds necessary. But we here assume human directional corrections $\boldsymbol{a}_{k_{i}}$ are in robot action space. For example, in autonomous driving where the robot action space is steering angle and acceleration, human can directly apply feedback with steering wheel and pedals. In some other applications, the action space correction can be achieved via specific mapping or interface. For example, to correct a torque-controlled robot arm, human’s correction on robot end-effector can be mapped into joint torque via Jacobian transpose matrices. The reason why we only consider the action space correction is because first, corrections in input spaces may be easier to implement, as shown in our later experiments, and second, corrections in state spaces can be infeasible for the under-actuated robot system [20].
After receiving each $\boldsymbol{a}_{k_{i}}$, the robot updates the safety parameter $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{i+1}$ in (1) for its MPC policy. Then, the robot continues with $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i+1}}^{\gamma}$ until the next possible user corrections. This process repeats until the convergence of the policy update or the user is satisfied with the robot motion. An illustration of one update is shown in Fig. 2.
Figure 2: Update of safe MPC policy with human directional corrections.
In the above learning process, we consider that a human user only provides directional correction $\boldsymbol{a}$, i.e., the magnitude $\left\lVert\boldsymbol{a}\right\rVert$ does not matter (theoretical justification will be given later). For example, when correcting a mobile robot, the user’s correction can be "left" or "right", without specifying how much the robot should go. As shown in our previous work [20] or later, the type of directional feedback provides a consistently large and flexible correction space, effectively overcoming the over-shooting correction, compared to existing magnitude feedback.
III-C3 Robot Operation Safety and Emergency Stop
In the aforementioned learning procedure, the robot will respond to human corrections by executing the updated safe MPC policy $\pi_{\boldsymbol{\theta}}^{\gamma}$. However, it is possible that the robot may still exit the user-intended safety region. In such cases, a human user may trigger an emergency stop mechanism to stop the robot motion and potentially restart it from a desired safety region. In this sense, all trajectories before the human triggers the emergency stop are considered safe in human intent. In other words, until an emergency stop is triggered, human corrections typically occur while the robot is still within the intended safety zone; otherwise, the corrections would be useless. For instance, correcting a vehicle after it is crashed is futile.
III-C4 Parametric Safety Constraints
In the safe MPC (4), the safety set is defined as a zero sublevel set of a safety value function ${g}_{\boldsymbol{\theta}}$. We note that if ${g}_{\boldsymbol{\theta}}$ has sufficient representation power, the safety set can represent complex (non-convex) safety region, which can be alternatively represented by the intersection of multiple constraints. In the following, we mainly focus on a weighted-features parameterization of ${g}_{\theta}$, which has been commonly used by the prior work on learning cost/reward [4, 20, 50, 51, 52]. Specifically, we parameterize $\boldsymbol{g}_{\theta}$
${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})=\phi_{0}(\boldsymbol{\xi})+\theta_{1}\phi_{1}(\boldsymbol{\xi})+\theta_{2}\phi_{2}(\boldsymbol{\xi})+\cdots+\theta_{r}\phi_{r}(\boldsymbol{\xi}),$ (5)
with $\boldsymbol{\theta}=[\theta_{1},\theta_{2},...\theta_{r}]^{\mkern-1.5mu\mathsf{T}}\in\mathbb{R}^{r}$ and $\boldsymbol{\phi}=[\phi_{1},\phi_{2},\dots,\phi_{r}]^{\mkern-1.5mu\mathsf{T}}$ the weight and feature vectors, respectively. We fix the weight of $\phi_{0}$ as unit, because otherwise, $\boldsymbol{\theta}$ has scaling ambiguity, i.e., any scaled $c\cdot\boldsymbol{\theta}$ with $c>0$ lead to the same MPC solutions. The dimension (number) $r$ of features can be chosen such that $\boldsymbol{\phi}$ has enough expressiveness to represent the human-intended safety constraint. We do not impose any constraints on the choice of feature vector as long as it is differentiable, for example, $\boldsymbol{\phi}$ can be neural features. In practice the feature vector can be chosen by trial and error, i.e., starting from a few features and gradually adding more until the learned constraint is satisfactory to the human user. Later, we will discuss our proposed method for generic parameterization $\boldsymbol{g}_{\boldsymbol{\theta}}$.
III-C5 Problem of Safe MPC Alignment
Following the procedure of human correction and robot policy update in Subsection III-C and given a weighted-feature safety constraint (5), we aim to develop a policy update rule: $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{i+1}$ for each received human directional correction, such that $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{H}\in\Theta_{H}$, as $i=1,2,3,\dots,$. i.e., safety constraints in the robot’s safe MPC policy will be aligned with human intent.
IV Overview of Safe MPC Alignment Algorithm
IV-A Hypothesis on Human Directional Corrections
Informally, we have the following intuitions on human feedback. First, human corrections, on average, improve the robot safety. Second, human correction on robot safety is typically proactive and happens before the robot leaves the human-intended safety set. In other words, human correction is made when the robot is still in $\mathcal{S}(\boldsymbol{\theta}_{H})$ in (4). For example, it is useless to correct a vehicle after its crash. Third, human feedback is typically safety-margin aware: as a robot approaches the safety boundary, human users tend to prioritize safety over the task objective. For instance, when a vehicle moves closer to obstacles, humans may instinctively shift their focus to avoiding collisions rather than pursuing the original target. In the following, we will formalize those intuitions.
Without loss of generality, suppose the motion plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ is a local solution to the penalty MPC (3), and when the robot execute the first action $\boldsymbol{u}_{0}^{\boldsymbol{\theta},\gamma}$ (receding horizon implementation), a human applies a directional correction $\boldsymbol{a}$ in robot action space. Define
$\mathbf{a}=(\boldsymbol{a}^{\mkern-1.5mu\mathsf{T}},0,0,...)\in\mathbb{R}^{mT}.$ (6)
We assume there is a nonzero volume set $\bar{\Theta}_{H}$ (containing a ball) in user implicit set $\Theta_{H}$, i.e., $\bar{\Theta}_{H}\subseteq\Theta_{H}$, such that for any $\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H}$, the human correction $\boldsymbol{a}$, on average, indicates a direction of lowering the penalty value $B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})$ for the robot motion plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$. Concretely, it means $\mathbf{a}$ on averages points in the direction of the negative gradient of $B(\boldsymbol{\xi},\boldsymbol{\theta}_{H})$ at $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, i.e.,
$\forall\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H},\ \operatorname{\mathbb{E}}_{\mathbf{a}}\left\langle\mathbf{a},-\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})\right\rangle=\\ \left\langle\mathbf{\bar{a}},-\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})\right\rangle\geq 0.$ (7)
[…]
(7) indicates that the magnitude of $\mathbf{a}$ does not affect the inequality. This brings one advantage of directional correction compared to the magnitude-based correction: regardless of robot plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, the allowable human corrections satisfying (7) always account for half of the action space, making directional corrections more accessible for human users, as shown in Fig. 3. Recall that $B(\boldsymbol{\xi},\boldsymbol{\theta})=J(\boldsymbol{\xi})-\gamma\ln(-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}))$, by extension of the gradient, (7) becomes
$\forall\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H},\left<\mathbf{\bar{a}},-\nabla J(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})\right>+\\ \gamma\frac{1}{-g_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})}\left<\mathbf{\bar{a}},-\nabla g_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})\right>\geq 0.$ (8)
Figure 3: Illustration of directional corrections in the space of robot motion plan $\boldsymbol{\xi}$s. The grey circles represent the level contours of $B(\boldsymbol{\xi},\boldsymbol{\theta}_{H})$. The black dot stands for current motion plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ and the red point stands for the plan $\boldsymbol{\xi}_{\boldsymbol{\theta}_{H}}^{\gamma}$ with human intended safety. The blue region stands for the half-space of all possible human corrections under our assumption. Some samples of human corrections are shown by blue arrows.
For the above equation, we have the following interpretations.
(I) The first and second terms on the left side of (8) correspond to human correction toward lowering the cost and improving the safety of $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, respectively. $\gamma$ serves as a weight between two improvements.
(II) If we interpret the value of $-{g_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})}$ as the safety margin of $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, (8) suggests that the human correction is safety-margin aware: the closer the robot motion get to the unsafe region, the more will human user emphasize on improving the robot’s safety than minimizing the objective. This property is consistent with our previous intuition.
(III) (7) implies $B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})$ is well-defined in the first place:
$\forall\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H},\ \quad g_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})<0.$ (9)
This corresponds to the third intuition: a user typically gives correction only when the robot is in user intended safety region, i.e., $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}\in\mathcal{S}(\boldsymbol{\theta}_{H})$.
In sum, we have used (7) and (9) to formalize the intuitions we have observed at the beginning of this session. Those two constraints on $\boldsymbol{\theta}_{H}$ will be used in the robot learning next.
IV-B Alignment Algorithm: Learning by Hypothesis Space Cutting
To find $\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H}\subseteq\Theta$ for the safe MPC (1), our idea is that for each human directional correction received $i=0,1,2,....$, the robot performs one learning iteration, by updating a hypothesis space $\Theta_{i}\in\mathbb{R}^{r}$, a set containing all possible $\bar{\Theta}_{H}$. At initial, the robot may choose a large enough hypothesis space $\Theta_{0}$. Specifically, one can choose $\Theta_{0}$ as a box space
$\Theta_{0}=\{\boldsymbol{\theta}\ |\ \boldsymbol{c}_{l}\leq\boldsymbol{\theta}\leq\boldsymbol{c}_{u}\}\subset\mathbb{R}^{r},$ (10)
in which $\boldsymbol{c}_{l},\boldsymbol{c}_{u}$ are the vectors of lower and upper bounds of the box space, resepctively, and the inequalities applied element-wise. By saying a $\Theta_{0}$ large enough, we assume
$\bar{\Theta}_{H}\subset\Theta_{0}.$ (11)
Later, we will show that our method can self-access the learning failure in case of misspecification, where (11) is not met.
Recall the procedure of human correction in Section III-C and Fig. 2. At $i$-th learning iteration, $i=1,2,3...$, the robot executes $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ with $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$. The user gives a directional correction $\boldsymbol{a}_{k_{i}}$ at rollout step ${k_{i}}$. By letting $\mathbf{\bar{a}}_{i}=(\boldsymbol{a}_{k_{i}}^{T},0,0,...)$ as in (6), we can establish the following two constraints on $\bar{\Theta}_{H}$ from the previous hypotheses (7) and (9):
$\bar{\Theta}_{H}\subset\mathcal{C}_{i}=\left\{\boldsymbol{\theta}|\left<\mathbf{\bar{a}}_{i},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta})\right>\leq 0,\,\,\,g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<0\right\},$ (12)
which we refer to as the cutting set. Note that the first inequality $\mathbf{\bar{a}}_{i},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta})\right>\leq 0$ is deduced by multiplying -1 on both sides of (7). Using $\mathcal{C}_{i}$, one can update the hypothesis space
$\Theta_{i}=\Theta_{i-1}\cap\mathcal{C}_{i}.$ (13)
The above update can be viewed as a step of using (12) to cut (intersect with) the current hypothesis space $\Theta_{i-1}$ to remove the portion where $\bar{\Theta}_{H}$ is impossible. The remaining hypothesis space $\Theta_{i}$ will be new hypothesis used for the next iteration. With the above update, we have the following result.
Lemma 2.
Given $\bar{\Theta}_{H}\subset\Theta_{0}$, if one follows the update (13), then $\bar{\Theta}_{H}\subset\Theta_{i}$ for all $i=1,2,3,...$.
Proof.
See the proof in Appendix A-A. ∎
As shown in Fig. 4, the procedure of robot learning is to cut and move hypothesis space based on human directional correction. Therefore, we name our alignment method Hypothesis Space Cutting. As a summary, we present the three main steps of our algorithm below.
Figure 4: The update of constraint parameter (2D examples). Two cutting lines are used to cut $\Theta_{i-1}$ to $\Theta_{i}$. The user implicit intent set $\bar{\Theta}_{H}$ is contained in $\Theta_{i}$. New $\boldsymbol{\theta}_{i+1}$ is chosen as the parameter for the next safe MPC policy.
Safe MPC alignment by Hypothesis Space Cutting Initialize $\Theta_{0}$ that contains $\bar{\Theta}_{H}$. Repeat the following three steps for $i=1,2,...,K$ until $\boldsymbol{\theta}_{i}\in\bar{\Theta}_{H}$. Step 1: Robot chooses parameter $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ for its current safe MPC policy $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$. Step 2: Human user applies a directional correction $\boldsymbol{a}_{k_{i}}$ while robot executes $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ Step 3: Perform hypothesis space cut using (13) to get new hypothesis space $\Theta_{i}$.
At $i$th learning iteration, the robot chooses $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ from the most recent hypothesis space $\Theta_{i-1}$ for its current MPC $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$. While executing $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, the robot receives human directional correction $\boldsymbol{a}_{k_{i}}$ at the rollout time step $k_{i}$. This correction leads to a cutting set $\mathcal{C}_{i}$ (12), which will be used to update $\Theta_{i-1}$ to $\Theta_{i}$ in (13). The above iteration repeats $i=1,2,3,...$. From (13), one can easily see that $\Theta_{0}\supseteq\Theta_{1}\supseteq...\supseteq\bar{\Theta}_{H}$. Given a volume measure, it follows $\operatorname{{Vol}}\Theta_{0}\geq\operatorname{{Vol}}\Theta_{1}\geq...,\geq\operatorname{{Vol}}{\bar{\Theta}_{H}}$. Thus, the volume of the hypothesis space is at least not increasing after each user correction. It is hoped that the volume of the hypothesis space decreases quickly, such that new $\boldsymbol{\theta}_{i+1}$ increases its chance of falling inside $\bar{\Theta}_{H}$ — a successful safety alignment between the robot and human.
For the above Safe MPC Alignment method, we will later (in Section V) show that within finite number $K$ of corrections, the method will either
  • (I)
    surely have $\boldsymbol{\theta}_{i}\in\bar{\Theta}_{H}\subset\Theta_{H}$, $i<K$, or
  • (II)
    declare the failure due to misspecification ($\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$)
Also in Section V, we will provide implementation details for the above hypothesis space cutting procedure.
IV-C Properties and Geometric Interpretations
In this session, we present some properties and the geometric interpretations for the proposed safe MPC alignment. Those properties will be used to establish the convergence of the algorithm in the next section.
IV-C1 Piecewise Linearity of Cuts
By recalling the weight-feature parameterization of the safety constraint (5), one can examine that the inequities in (12) are both linear in $\theta$. This means that the cutting set $\mathcal{C}_{i}$ in (12) is a piecewise linear cut, as stated formally in the following lemma.
Lemma 3.
$\forall i$, $\mathcal{C}_{i}$ in (12) can be expressed as an intersection of two half-spaces in $\mathbb{R}^{r}$:
$\mathcal{C}_{i}=\{\boldsymbol{\theta}\ |\ \boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}\leq b_{i}\}\cap\{\boldsymbol{\theta}\ |\ {\boldsymbol{\theta}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<-\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\},$ (14)
with
$\small\boldsymbol{h}_{i}=-\left<\mathbf{\bar{a}}_{i},\nabla J(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>\boldsymbol{\phi}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})+\gamma\frac{\partial\boldsymbol{\phi}}{\partial\boldsymbol{u}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\mathbf{\bar{a}}_{i},$ (15)
[…]
Proof.
See Appendix A-B. ∎
According to the above lemma, $\Theta_{i-1}\cap\mathcal{C}_{i}$ can be viewed as cutting $\Theta_{i-1}$ with the boundary of $\mathcal{C}_{i}$ which is two hyperplanes:
$\begin{split}\mathbf{bd}\ \mathcal{C}_{i}=\{\boldsymbol{\theta}\ |\ \boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}=b_{i},{\boldsymbol{\theta}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\leq-\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\}\ \cup\\ \{\boldsymbol{\theta}\ |\ \boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}\leq b_{i},{\boldsymbol{\theta}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})=-\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\},\end{split}$ (17)
shown in Fig. 4.
As the initial hypothesis space $\Theta_{0}\in\mathbb{R}^{r}$ in (10) is a convex polytope, iteratively cutting with hyperplane pairs yields a series of convex polytopes in $\mathbb{R}^{r}$. Consequently, all $\Theta_{i}$s, $i=0,1,2,...$, are convex polytopes.
IV-C2 Relationship between $\boldsymbol{\theta}_{i}$ and $\mathbf{bd}\ \mathcal{C}_{i}$
According to (14), the cutting set $\mathcal{C}_{i}$ depends on plan $\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ which is a solution to penalty-based MPC (3) with $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$. The following assertion states a relationship between $\boldsymbol{\theta}_{i}$ and $\mathbf{bd}\ \mathcal{C}_{i}$.
Lemma 4.
$\forall i,\ \boldsymbol{\theta}_{i}$ is on $\mathbf{bd}\ \mathcal{C}_{i}$ in (17). Specifically, $\boldsymbol{\theta}_{i}$ satisfies
$\boldsymbol{\theta}_{i}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}=b_{i}\quad\text{and}\quad\boldsymbol{\theta}_{i}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<-\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})$ (18)
for all $i=1,2,...$.
Proof.
See Appendix A-C. ∎
Lemma 4 states $\boldsymbol{\theta}_{i}\in\mathbf{bd}\ \mathcal{C}_{i}$. Its geometric interpretation is in Fig. 5: the cut $\mathbf{bd}\ \mathcal{C}_{i}$ always passes $\boldsymbol{\theta}_{i}$. This suggests that the choice of $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ is critical in the learning process: it directly affects the position of the cut $\mathbf{bd}\ \mathcal{C}_{i}$ and consequently the removal volume. It is important to choose a good $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ such that a large hypothesis space is removed after each cut.
We briefly discuss the implication of different choices of $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$. When choosing $\boldsymbol{\theta}_{i}$ near the boundary of $\Theta_{i-1}$, as shown in Fig. 5 and Fig. 5, the removal of the hypothesis space depends on the specific direction of cut $\mathbf{bd}\,\mathcal{C}_{i}$. Since the direction of the cut $\mathbf{bd}\,\mathcal{C}_{i}$ will be eventually determined by the direction of user correction according to (17), the removal volume of the hypothesis space can be relatively small, as in Fig. 5, or large, as in Fig. 5, depending on the specific direction of the correction.
As the specific direction of user correction is hard to assume, one can intuitively see that a good choice of $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ will be near the center of $\Theta_{i-1}$. As such, it can lead to an averaged large cut of the hypothesis space regardless of the direction of $\mathbf{bd}\ \mathcal{C}_{i}$, as shown in Fig. 5. We will rigorously discuss the centering choice of $\boldsymbol{\theta}_{i}$ and the guarantee of volume reduction rate of the hypothesis space in the next section.
Figure 5: Illustration of the different choices of $\boldsymbol{\theta}_{i}$. (a) and (b): When $\boldsymbol{\theta}_{i}$ is chosen near the set boundary, the removal volume depends on the specific direction of $\mathbf{bd}\ \mathcal{C}_{i}$. (c): When $\boldsymbol{\theta}_{i}$ is near the center, the cut volume is large on average regardless of the specific user directional correction.
V Main Results and Implementation Details
In this section, we will present the details for how to select $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ at each learning iteration $i$ in our Safe MPC Alignment algorithm. Based on the previous geometric intuition, our choice of $\boldsymbol{\theta}_{i}$ is to achieve a large averaged reduction rate in the hypothesis space volume for all possible human directional corrections. The reduction rate enables us to establish an exponential convergence of the learning process. As a result, we can find an upper bound of total human corrections for successful learning, i.e., $\boldsymbol{\theta}_{i}\in\bar{\Theta}_{H}$, or identify the misspecification of hypothesis space.
V-A The Choice of $\boldsymbol{\theta}_{i}$
As discussed previously, $\boldsymbol{\theta}_{i}$ needs to be picked from the "center" of $\Theta_{i-1}$ in each learning iteration. One option is to choose the mass center of $\Theta_{i-1}$. However, computing the centroid of a polytope can be #P-Hard [53]. While methods like the Hit-And-Run algorithm [54] estimate the mass center by sampling, they can suffer from high sample complexity. Another possible choice is to use the center of the maximum inscribed ball of a polytope [55], but it lacks of the guarantee in volume reduction rate. For computational efficiency and guaranteed convergence, we will choose $\boldsymbol{\theta}_{i}$ as the center of the Maximum Volume Ellipsoid (MVE) inscribed in the convex set $\Theta_{i-1}$, shown in Fig. 6.
Figure 6: $\boldsymbol{\theta}_{i}$ is chosen as the center of MVE of $\Theta_{i-1}$ (2D example).
Definition 1 (Maximum Volume inscribed Ellipsoid (MVE)[56]).
The MVE of a given convex set $\Theta\subset\mathbb{R}^{r}$ is defined as the image of the unit ball under an affine transformation
=$\mathcal{E}=\{\bar{H}\boldsymbol{\theta}+\bar{\boldsymbol{d}}\ |\ \|\boldsymbol{\theta}\|_{2}\leq 1\},$ (19)
where the affine transformation parameter $\bar{H}\in\mathbb{S}^{r}_{++}$ (a $r\times r$ symmetric positive definite matrix), and $\bar{\boldsymbol{d}}\in\mathbb{R}^{r}$ is called the center of $\mathcal{E}$. $\bar{H},\bar{\boldsymbol{d}}$ are determined via the following optimization problem to maximize the volume of the ellipsoid $\det H$:
$\begin{split}\bar{H},\bar{\boldsymbol{d}}=\underset{H\in\mathbb{S}^{r}_{++},\boldsymbol{d}}{\arg\max}\quad&\log\det H\\ \rm{s.t.}\quad&\sup\nolimits_{\|\boldsymbol{\theta}\|_{2}\leq 1}\mathbbm{1}_{\Theta}(H\boldsymbol{\theta}+\boldsymbol{d})\leq 0,\end{split}$ (20)
where the indicator function $\mathbbm{1}_{\Theta}(\boldsymbol{\theta})=0$ for $\boldsymbol{\theta}\in\Theta$ and otherwise $\mathbbm{1}_{\Theta}(\boldsymbol{\theta})=+\infty$. Specifically, when $\Theta$ is a polytope $\{\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{n}_{i}\leq c_{i},\ i=1,\dots,m\}$, the problem can be converted to a convex optimization with $m$ constraints, see chapter 8.4.2 in [56]:
$\begin{split}\bar{H},\bar{\boldsymbol{d}}=\underset{H\in\mathbb{S}^{r}_{++},\boldsymbol{d}}{\arg\min}\quad&\log\det H^{-1}\\ \rm{s.t.}\quad&\left\lVert H\boldsymbol{n}_{i}\right\rVert_{2}+\boldsymbol{d}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{n}_{i}\leq c_{i},\ i=1,...,m.\end{split}$ (21)
Recall that in our Safe MPC Alignment the learning update (13) is to iteratively cut the hypothesis space by a pair of hyperplanes (17), and that the initial hypothesis space (11) is a box. By induction, it leads to that any updated hypothesis space $\Theta_{i}$ is a convex polytope, $i=0,1,2,...$. Thus, for any $\Theta_{i-1}$, by the above definition, let $\mathcal{E}_{i-1}$ denote the MVE of $\Theta_{i-1}$ and $\bar{\boldsymbol{d}}_{i-1}$ denote the center of $\mathcal{E}_{i-1}$. Then, the choice of $\boldsymbol{\theta}_{i}$ is
$\boldsymbol{\theta}_{i}=\bar{\boldsymbol{d}}_{i-1}.$ (22)
The convex optimization (20) can be solved efficiently [57].
V-B Result 1: Finite Iteration Convergence
When choosing $\boldsymbol{\theta}_{i}$ as the center of MVE of $\Theta_{i-1}$, the following lemma states that the volume ratio ${\mathbf{Vol}(\Theta_{i})}/{\mathbf{Vol}(\Theta_{i-1})}$ after each hypothesis space cut can be upper bounded.
Lemma 5.
In the Safe MPC Alignment algorithm, when choosing $\boldsymbol{\theta}_{i}\in\mathbb{R}^{r}$ as the center of MVE of $\Theta_{i-1}$, one has
$\frac{\mathbf{Vol}(\Theta_{i})}{\mathbf{Vol}(\Theta_{i-1})}\leq 1-\frac{1}{r},$ (23)
for all $i=1,2,3,...$ and any enclosure volume measure $\bf{Vol}$.
Proof.
See Appendix A-D. ∎
The lemma above asserts an upper bound of the volume of the updated hypothesis space $\mathbf{Vol}(\Theta_{i})$ relative to the previous volume. With this, we can now establish the finite iteration convergence of the Safe MPC Alignment, stated below.
Theorem 1 (Successful learning in finite iterations).
In the Safe MPC Alignment algorithm, $\boldsymbol{\theta}_{i}$ is selected as the center of MVE of $\Theta_{i-1}$, $i=1,2,3,...$. If $\bar{\Theta}_{H}\subset\Theta_{0}$, the algorithm will terminate successfully within
$K=\left\lceil\frac{\ln\frac{\tau_{r}\rho_{H}^{r}}{\mathbf{Vol}(\Theta_{0})}}{\ln(1-\frac{1}{r})}\right\rceil.$ (24)
iterations, where $\rho_{H}$ is the radius of a ball contained within $\bar{\Theta}_{H}$ and $\tau_{r}$ is the volume of a unit ball in $\mathbb{R}^{r}$ space. That is, there exists $i<K$, such that $\boldsymbol{\theta}_{i}\in\bar{\Theta}_{H}\subset\Theta_{H}$.
Proof.
See Appendix A-E. ∎
The above main result asserts a maximum iteration count for the proposed algorithm to find $\boldsymbol{\theta}_{i}\in\Theta_{H}$. In other words, within $K$ rounds of human corrections, the robot can successfully learn a safety constraint intended by the user. The above result requires a pre-set of $\rho_{H}^{r}$, which is the radius of a ball contained in $\bar{\Theta}_{H}$. Since the size of $\bar{\Theta}_{H}$ is unknown in practice, $\rho_{H}>0$ can be interpreted as an underestimated size of the unknown user intent set $\bar{\Theta}_{H}$. It also serves as a termination threshold for learning convergence: a smaller $\rho_{H}$ can lead to a more concentrated of $\boldsymbol{\theta}_{H}$ learned at convergence.
However, the finite iteration convergence in Theorem 1 comes with a prerequisite: $\bar{\Theta}_{H}\in\Theta_{0}$ has to be satisfied in the first place in the Safe MPC Alignment algorithm. Although such a condition can be achieved by empirically choosing $\Theta_{0}$ large enough to contain all possible $\bar{\Theta}_{H}$ (10), it is also desired to have a guaranteed method to detect the hypothesis misspecification, where such condition is not met. Next, we will discuss the declaration of hypothesis misspecification.
V-C Result 2: Hypothesis Misspecification Declaration
We consider the case of hypothesis space misspecification, where $\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$. In this case, no $\boldsymbol{\theta}\in\Theta_{0}$ can account for the user intended safety (4). As shown in [52, 58], the hypothesis misspecification can lead to the undesired behavior of the robot. The following result states the proposed Safe MPC Alignment algorithm can automatically certify the misspecification of the initial hypothesis space.
Theorem 2 (Certifying misspecification).
In the Safe MPC Alignment algorithm, $\boldsymbol{\theta}_{i}$ is selected as the center of MVE of $\Theta_{i-1}$, $i=1,2,3,...$. If $\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$, the algorithm converges to a set $\Theta^{*}$ as $i\rightarrow\infty$, such that $\mathbf{Vol}(\Theta^{*})=0$ and $\Theta^{*}\cap\mathbf{int}\ \Theta_{0}=\emptyset$.
Proof.
See Appendix A-F. ∎
The above result states that when the initial hypothesis space is misspecified, $\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$, the Safe MPC Alignment algorithm will finally converge to a set of zero volume on the boundary of the initial hypothesis space $\Theta_{0}$.
The above theorem inspires us to empirically identify the misspecification by monitoring the position of the MVE center of $\Theta_{i}$ during the learning update. As shown in Fig. 7, if after the maximum iteration $K-1$ in (24), the MVE center $\boldsymbol{\theta}_{K-1}$ is very close to the boundary of the initial hypothesis set $\Theta_{0}$, there is likely hypothesis space misspecification, and the algorithm terminates with misspecification failure.
Figure 7: When there is misspecification, the Safe MPC Alignment algorithm will converge to some point on the boundary of $\Theta_{0}$. So we can use the distance of $\boldsymbol{\theta}_{K-1}$ to the boundary as an indicator for misspecification.
Specifically, recall that $\Theta_{0}$ in (10) is a box with bounds $\boldsymbol{c}_{l},\boldsymbol{c}_{h}$, we can pre-define a small threshold $\epsilon$ to examine if $\boldsymbol{\theta}_{K-1}$ is on the boundary of $\Theta_{0}$. If $\boldsymbol{\theta}_{K-1}$ satisfies
$\min_{j{=}1,...,r}\ \min\big{(}\lvert\boldsymbol{c}_{l}[j]-\boldsymbol{\theta}_{K-1}[j]\rvert,\lvert\boldsymbol{c}_{u}[j]-\boldsymbol{\theta}_{K-1}[j]\rvert\big{)}\leq\epsilon$ (25)
it means $\boldsymbol{\theta}_{K-1}$ is very close to the boundary of $\Theta_{0}$. We consider there is likely misspecification in this case.
V-D Detailed Algorithm of Safe MPC Alignment
With the main results on convergence and misspecification, we can establish the full safe MPC alignment algorithm in Algorithm 1. With initializing $\Theta_{0}$ and pre-define a termination threshold $\rho_{H}$, three steps in the hypothesis space cutting are executed until the user satisfies robot motion or the iteration count reaches the maximum count $K$ in (24). Note that during the execution of the MPC policy, $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, the human can perform emergency stopping and reset the robot’s state, which corresponds to Section III-C3. After $K-1$ iterations, the algorithm checks the misspecification by (25).
Initialization: Parameterized constraint $\boldsymbol{g}_{\theta}$, initial hypothesis space $\Theta_{0}$, termination threshold $\rho_{H}$ (i.e., approximate radius of $\bar{\Theta}_{H}$), $i=1$
Compute maximum correction count $K$ from $\rho_{H}$ in (24);
while $i<K$ do
       Choose $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ using (22);
      Robot executes $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$;
      if Human provides directional correction $\boldsymbol{a}_{k_{i}}$ then
             Generate $\bar{\boldsymbol{\mathbf{a}}}_{i}$ using (6);
             Calculate the cutting set $\mathcal{C}_{i}$ using (14);
             Update hypothesis space $\Theta_{i}=\Theta_{i-1}\cap\mathcal{C}_{i}$;
             $i\leftarrow i+1$;
      else if  Human triggers emergency stop  then
             Human restarts robot from safe region.;
      else if  Human user is satisfied  then
             return $\boldsymbol{\theta}_{i}$;
end while
if $\boldsymbol{\theta}_{K-1}$ satisfies (25) then
       return misspecification failure;
else
       return $\boldsymbol{\theta}_{K-1}$
end if
Algorithm 1 Safe MPC Alignment Algorithm
VI Numerical Examples and Analysis
In this section, we will evaluate the proposed Safe MPC Alignment using two numerical examples: inverted pendulum and two-link torque-controlled robot arm moving in constrained environment. Simulated human corrections knowing ground-truth safety constraints will be used. We will compare the proposed method with two baselines, and conduct ablation studies on the hyperparameters of the algorithm. All numerical experiments are implemented in Python, running on a Dell desktop with Intel i9-13900K CPU. All our experiment codes are open-sourced at at https://github.com/asu-iris/Safe-MPC-Alignment.
VI-A Inverted Pendulum
VI-A1 Setup
The dynamics of an inverted pendulum is
$\ddot{\alpha}=\frac{3}{ml^{2}}(-\frac{1}{2}m\mathrm{g}l\sin{\alpha}+u-d\dot{\alpha}),$ (26)
with $\alpha$ being the angle between the pendulum and gravity, $\mathrm{g}$ the gravity constant, $u$ input torque, $m$, $l$ and $d$ the length, inertia, and damping ratio. We set the system state as $\boldsymbol{x}=[\alpha,\dot{\alpha}]^{\mkern-1.5mu\mathsf{T}}$ and discretize (26) by the Euler method with $\triangle t=0.02\mathrm{s}$. A zero-mean Gaussian noise with covariance $\Sigma=\mathbf{diag}(1,4)\times 10^{-5}$ is added to $\boldsymbol{x}$ at each control step to simulate the state noise. The task is to control the pendulum to the target $\boldsymbol{x}^{*}=[\pi,0]^{\mkern-1.5mu\mathsf{T}}$. We additionally pose an unknown global state constraint on pendulum motion:
$0.6\alpha+\dot{\alpha}=\boldsymbol{\theta}^{H}\boldsymbol{x}\leq 3,$ (27)
where $\boldsymbol{\theta}^{H}=[0.6,1]^{\mkern-1.5mu\mathsf{T}}$ is to be learned. All other parameter values are given in Appendix.
In Safe MPC (1), we set $T=40$ and the cost functions as
$\begin{gathered}c(\boldsymbol{x}_{t},u_{t})=Ru_{t}^{2}\\ h(\boldsymbol{x}_{T})=(\boldsymbol{x}_{T}-\boldsymbol{x}^{*})^{T}Q(\boldsymbol{x}_{T}-\boldsymbol{x}^{*}),\end{gathered}$ (28)
respectively, with $R=0.1$ and $Q=\mathrm{diag}(25,10)$. To ensure the safe MPC policy to satisfy a global constraint, it is sufficient to learn a safety constraint defined on the closest controllable state. Thus, in (1), we define
${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})=-3+\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{x}_{1}.$ (29)
In the penalty-based MPC (3), we set $\gamma=0.1$.
Suppose the user intent set $\bar{\Theta}_{H}=\{\boldsymbol{\theta}|\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert\leq 0.02\}$. The simulated human corrections $\mathbf{a}$ are generated as follows: when the system state is near the safety boundary, i.e., ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})>-\epsilon_{g}$ ($\epsilon_{g}=0.25$ here), a directional correction $\mathbf{a}$ is be generated in probability $p=0.3$ using $\mathrm{sign}(-\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H}))$ (from which only the first entry is taken considering (6)). The correction is probabilistic to simulate the randomness of human providing corrections. It is easy to verify that the simulated $\mathbf{a}$ is valid in term of satisfying the hypotheses in section IV-A.
VI-A2 Convergence Analysis
We first evaluate the convergence of our algorithm without hypothesis misspecification. In Algorithm 1, we set the initial hypothesis $\Theta_{0}$ with box boundaries $\boldsymbol{c}_{l}=[-6,-6]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{u}=[2,2]^{\mkern-1.5mu\mathsf{T}}$; thus $\Theta_{0}$ includes $\bar{\Theta}_{H}$. The initial $\boldsymbol{\theta}_{1}=[-2,-2]^{\mkern-1.5mu\mathsf{T}}$ (i.e., the center of $\Theta_{0}$).
With each simulated human correction, we track $\boldsymbol{\theta}_{i}$ in each learning update. Each update of ${\Theta}_{i}$ takes less than 30ms. $\boldsymbol{\theta}_{i}$ falls in the ground-truth set $\bar{\Theta}_{H}$ after 14 corrections. The MPC rollout trajectory (phase plot) with each $\boldsymbol{\theta}_{i}$ is shown in Fig. 8. The ground-truth constraint (27) and the learned constraints $\boldsymbol{g}_{\boldsymbol{\theta}_{i}}$ are also plotted in Fig. 8. From Fig. 8, we can conclude that the ground-truth constraint can be successfully learned from simulated corrections.
(a) Initial trajectory (with $\boldsymbol{\theta}_{1}$)
(b) After 3 corrections (with $\boldsymbol{\theta}_{4}$)
(c) After 6 corrections (with $\boldsymbol{\theta}_{7}$)
(d) After 14 corrections (with $\boldsymbol{\theta}_{15}$)
Figure 8: The pendulum rollout trajectory at different correction interactions: (a) trajectory with $\boldsymbol{\theta}_{1}$, (b) with $\boldsymbol{\theta}_{4}$, (c) with $\boldsymbol{\theta}_{7}$, (d) with $\boldsymbol{\theta}_{15}$. The learned constraint with $\boldsymbol{\theta}_{1}$ is omitted since it covers the whole first quadrant. As the learning proceeds, the learned constraint boundary (orange line) converges to the ground-truth constraint boundary (red line). When the learning process terminates, the learned and ground-truth constraint boundaries almost overlap, suggesting the success of constraint learning from corrections.
To further analyze the learning convergence, we run the algorithm for 10 times. In each run, we keep track of two metrics over correction (iteration) count and show them in Fig. 9. The first metric (left panel) is the logarithm volume measure of the solved MVE $\log\det\bar{H}$ in (20); this provides a lower bound of $\Theta_{i}$ to show the exponential convergence property. The second (right panel) is the parameter error $\left\lVert\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{H}\right\rVert_{2}$. From Fig. 9, we can conclude the proposed method is
  • (I)
    Fast-converging. The logarithm volume $\log\det\bar{H}$ has near-linear decreasing speed. This is consistent with the exponential convergence in theory stated in Lemma 5.
  • (II)
    Correction-efficient. $\left\lVert\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{H}\right\rVert_{2}$ converges to the value lower than 0.02 within 17 corrections. This is consistent with the theoretical upper bound of correction count from given by Theorem 1 :
    ${\ln\frac{\tau_{r}\rho_{H}^{r}}{\mathbf{Vol}(\Theta_{0})}}/{\ln(1-\frac{1}{r})}=17.28.$ (30)
(a) Logarithm MVE Volume
(b) Parameter Estimation Error
Figure 9: (a) The plot of $\log\det\bar{H}$ with respect to correction numbers. In all algorithm runs, the plot shows the near-linear decrease of the logarithm volume, suggesting the exponential decrease of the volume of MVE of the hypothesis spaces, as asserted in Theorem 1. (b) The plot of the parameter estimation error $\left\lVert\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{H}\right\rVert_{2}$ versus correction numbers. In all runs, the parameter estimation error can converge to a value lower than 0.02 within 17 corrections.
VI-A3 Analysis for Misspecification Case
To analyze the performance of our algorithm to certify misspecification cases, we set the initial hypothesis space $\Theta_{0}$ as ${\boldsymbol{c}_{l}}=[-1,-1]^{\mkern-1.5mu\mathsf{T}}$ and ${\boldsymbol{c}_{u}}=[0.8,0.8]^{\mkern-1.5mu\mathsf{T}}$ which does not contain $\bar{\Theta}_{H}$. In this setting, the algorithm automatically stops after 7th correction. The path of $\boldsymbol{\theta}_{i},\ i=1,\dots,8$ is shown in Fig. 10. In Fig. 10, $\boldsymbol{\theta}_{i}$ is continuously moving towards the boundary of $\Theta_{0}$. This fact verifies Theorem 2 and justifies the proposed method is capable of misspecification detection.
Figure 10: Illustration of misspecification case. The blue dashed lines are the boundary of $\Theta_{0}$. $\boldsymbol{\theta}_{i},\ i=1,\dots,7$ is shown in he orange dots in one run. Black arrows show the update directions of $\boldsymbol{\theta}_{i}$s. The center of the ground-truth intent set is marked as the red star, and the $\boldsymbol{\theta}_{8}$ at the termination of learning is marked as the green triangle. When misspecification happens, $\boldsymbol{\theta}$ will converge toward the boundary of $\Theta_{0}$, as stated by Theorem 2.
VI-B Two-link Robot Arm Reaching with Complex Constraint
VI-B1 Setup
The dynamics if torque controlled two-link robot arm moving in the horizontal plane is $M(\boldsymbol{q})\ddot{\boldsymbol{q}}+\boldsymbol{c}(\boldsymbol{q},\dot{\boldsymbol{q}})=\boldsymbol{\tau}$, where $M(\boldsymbol{q})$ is the inertia matrix; $\boldsymbol{c}(\boldsymbol{q},\dot{\boldsymbol{q}})$ is the Coriolis term. $\boldsymbol{q}=[q_{1},q_{2}]^{\mkern-1.5mu\mathsf{T}}$ defines the two joint angles of the robot. The system state is defined as $\boldsymbol{x}=[\boldsymbol{q},\dot{\boldsymbol{q}}]$. The control input $\boldsymbol{u}$ is the joint torque $\boldsymbol{\tau}=[\tau_{1},\tau_{2}]^{\mkern-1.5mu\mathsf{T}}$, i.e. $\boldsymbol{u}=\boldsymbol{\tau}$. The other detailed physical parameters are given in Appendix.
As shown in Fig. 12, the robot starts at $\boldsymbol{x}_{\mathrm{init}}=[-\frac{\pi}{2},0,0,0]^{\mkern-1.5mu\mathsf{T}}$ and the task is to reach a target $\boldsymbol{x}^{*}=[\frac{\pi}{2},0,0,0]^{\mkern-1.5mu\mathsf{T}}$. In Safe MPC (1), we set $T=30$, the step cost functions
$\begin{gathered}c(\boldsymbol{x}_{t},\boldsymbol{u}_{t})=\boldsymbol{x}_{t}^{\mkern-1.5mu\mathsf{T}}P\boldsymbol{x}_{t}+\boldsymbol{u}_{t}^{\mkern-1.5mu\mathsf{T}}R\boldsymbol{u}_{t},\\ h(\boldsymbol{x}_{T})=(\boldsymbol{x}_{T}-\boldsymbol{x}^{*})^{T}Q(\boldsymbol{x}_{T}-\boldsymbol{x}^{*}),\end{gathered}$ (31)
with $P=\mathrm{diag}(1,1,2,2)$, $R=\mathrm{diag}(0.1,0.1)$ and $Q=\mathrm{diag}(5,5,10,10)$.
As shown by the black lines (obstacles) in Fig. 12, we define a complex reaching environment with non-convex obstacles and narrow corridor. The ground-truth global safety constraint corresponding to Fig. 12 is given
${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi})=-1+\boldsymbol{\theta}_{H}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}$ (32)
with $\boldsymbol{\phi}{=}[\mathrm{tanh}(x_{e}y_{e}),\mathrm{tanh}(x_{e}^{2},\mathrm{tanh}(y_{e}^{2}),\mathrm{tanh}(x_{e}),\mathrm{tanh}(y_{e})]^{\mkern-1.5mu\mathsf{T}}$, $\boldsymbol{\theta}_{H}=[0.78,0.91,-0.52,0.52,-1.04]^{\mkern-1.5mu\mathsf{T}}$, and $[x_{e},y_{e}]^{\mkern-1.5mu\mathsf{T}}$ is the end-effector position of the robot arm. The human intent set is $\bar{\Theta}_{H}{=}\{\boldsymbol{\theta}|\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert\leq 0.2\}$. We set the initial hypothesis $\Theta_{0}$ with box boundaries $\boldsymbol{c}_{l}=[-3,\dots,-3]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{u}=[3,\dots,3]^{\mkern-1.5mu\mathsf{T}}$. The simulated human corrections are generated in the similar way to that in Section VI-A.
Figure 11: The heatmap of a learned constraint $g$. The learned and ground-truth constraints are almost identical, showing the success of our method for learning complex constraint from corrections.
VI-B2 Result
Our method can successfully learn the safety constraint within 25-30 simulated corrections. One instance of the learned constraint is shown in Fig. 11. The comparison of the robot reaching motion without and with the learned safety constraint is shown in Fig. 12. The robot is capable of operating in the intended safe region after the algorithm converges. Those results show that the efficacy of the proposed method for solving system with complex safety constraints.
(a) Trajectory Before Corrections
(b) Trajectory After Corrections
Figure 12: The trajectory of the two-link robot arm (a) without and (b) with learned safety constraint. The robot successfully achieves the task with the learned safety constraint.
VI-C Baseline Comparison
In this session, we compare our method with two baseline methods using magnitude corrections in the pendulum environment (the task and model settings are identical to Section VI-A). As reviewed previously, the direct literature learning safety constraints from human feedback is sparse to the authors’ best knowledge. Thus, the chosen baselines are an adaptation from the most relevant methods.
Baseline 1: gradient matching. This baseline is inspired by [59], where raw human correction (with magnitude) is directly used as the gradient direction. In a similar spirit, when a human feedback $\mathbf{a}$ is made on robot trajectory $\xi_{\boldsymbol{\theta}}^{\gamma}$, we directly use the human magnitude correction as the true gradient label to learn the constraint $\boldsymbol{\theta}$. This leads to a learning loss
$\mathcal{L}=\left\lVert\mathbf{a}+\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})\right\rVert^{2}$ (33)
Baseline 2: maximum likelihood learning. This baseline follows [60], where it views $e^{{g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})}$ as a probability of human applying the "stop-feedback" to robot motion $\boldsymbol{\xi}$, when the robot reaches in a $\epsilon_{g}$-proximity of the safety boundary, i.e., when ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})>-\epsilon_{g}$ Thus, a dataset of trajectory-correction paris, i.e., $\mathcal{D}=\{\mathbf{a}_{i},\boldsymbol{\xi}_{i}\}_{i=0}^{N}$, can be used as the training data to learn $\boldsymbol{\theta}$ using maximum likelihood, while also penalizing positivity of ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{i})$ as a regulizer term. This leads to the following loss
$\mathcal{L}=\sum\nolimits_{i=1}^{N}\big{(}-{g}_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{i})+\alpha\cdot\mathrm{softplus}(g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{i}))\big{)}$ (34)
In our comparison, we choose $N=15$ (similar to correction number needed for our method to converge) and $\alpha=5$ for its best performance.
The comparison between our method and Baseline 1 focuses on data efficiency. The convergence criteria of $\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert\leq 0.02$ is set for both methods. 50 independent runs are performed and the number of correction needed for convergence are counted. We use Adam optimizer [61] with learning rate 0.02. We summarize the results in Table I. The parameter estimation error $\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert$ versus learning iteration (correction count) for a best run (i.e., best performance for the baseline) is plotted in Fig. 13(a). The comparison results demonstrate the significant data efficiency of the proposed method compared to gradient matching baseline, even though the latter uses both magnitude and direction information of corrections to learn.
TABLE I: Comparison of Safe MPC Alignment and Baseline I
Method Correction count for convergence
Safe MPC Alignment $11.76\pm 2.14$
Gradient Matching $240.38\pm 120.70$
(a) Comparison with Baseline 1
(b) Comparison with Baseline 2
Figure 13: (a) Comparison with gradient matching baseline: the parameter estimation error $\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert$ versus learning iteration (correction count). This is best run (among 50 runs) with the least corrections for the baseline. Our method converges after 12 corrections, while the baseline needs more than 160 corrections to converge. (b) Comparison with the maximum likelihood baseline under the different $\epsilon_{g}$. In the plot, the y-axis is $\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert$ at convergence and the error bar denotes the standard deviation. The maximum likelihood baseline fails to learn the correct weights when $\epsilon_{g}$ is large, while our method consistently shows high learning accuracy.
The comparison between our method and maximum likelihood baseline focuses on the varying positions of simulated correction, i.e., varying $\epsilon_{g}$. Recall that the simulated corrections will be generated only when ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})>-\epsilon_{g}$. The value of $\epsilon_{g}$ determines the position of the correction applied. In our comparison, we vary $\epsilon_{g}$ from 0.1 to 0.9 and track the converged parameter error for both methods. As shown in Fig. 13(b), safe MPC alignment consistently converges to the ground-truth $\boldsymbol{\theta}_{H}$ regardless of $\epsilon_{g}$, while the maximum likelihood baseline fails when $|\epsilon_{g}|$ is large. The results suggest the robustness of the proposed method against different correcting conditions compared to maximum likelihood baseline method.
VI-D Ablation Study
In this session, we conduct two ablation studies to show the performance of our algorithm given different design parameters/aspects. The task is still based on pendulum environment with settings identical to Section VI-A).
VI-D1 Choosing $\boldsymbol{\theta}_{i}$: MVE centering versus random sampling
In the first study, we compare two strategies in selecting $\boldsymbol{\theta}_{i}$ from the hypothesis space $\Theta_{i-1}$ in our algorithm: MVE centering (used in our method) and random sampling. For the latter, $\boldsymbol{\theta}_{i}$ is chosen by sampling from the uniform distribution supported by $\Theta_{i-1}$. We compare the total number of corrections needed at learning convergence under the two strategies.
With 50 independent runs, we observe that the MVE centering requires $11.76\pm 2.14$ corrections, while random sampling requires $11.88\pm 3.50$ corrections. While the mean value of the correction count looks similar, the random sampling shows a slightly larger variance. This is understandable given the sampling randomness. We also observed that sampling process introduced large computational overhead, which is around 25% time slower than MVE centering. In contrast, the MVE centering used in our algorithm shows low variance and computational efficiency in finding $\boldsymbol{\theta}_{i}$. Because selecting $\boldsymbol{\theta}_{i}$ is casted as a convex programming for MVE centering, such advantage would be more significant for high-dimensional parameter space.
VI-D2 Choosing $\gamma$
In the second study, we investigate how the setting of $\gamma$ affects the learning performance. We vary $\gamma$ from 0.01 to 10. Fig. 14 shows the total number of corrections needed at convergence versus different $\gamma$s.
The result in Fig. 14 shows that the learning performance remains the same given different settings of $\gamma$. Since $\gamma$ is weights parameter to control the safety conservativeness in the penalty-based MPC policy in (3), it means our method can successfully learn a safety constraint when the robot executes with safe MPC policy of different conservativeness levels.
Figure 14: The total number of corrections needed for convergence under different $\gamma$, the x-axis is of log scale. Safe MPC alignment can converge with a similar number of corrections (iterations) under different levels of $\gamma$.
VII User Study in Simulation Games
In this section, we test the Safe MPC Alignment method in two real-human user studies using two simulation games developed based on MuJoCo [62]. One is the drone navigation game and the other the Franka robot arm reaching game. In both games, the robot is controlled by an MPC controller and human participants provide real-time corrections through a keyboard. The objective of each game is to teach a robot to successfully learn a safe MPC to move through environments with unknown obstacles. In following, Subsection VII-A describes the basic setting of the two games. Subsection VII-B details the user study procedure and the recruitment of human participants. The results and analysis are given in VII-C.
VII-A Setting of Two Simulation Games
In both user studies, human participants are asked to give correction to the following two physics-based simulation games developed based on MuJoCo.
(1) Drone Navigation Game. As shown in Fig. 15, the goal of this game is to let a drone to successfully navigate through a narrow gate (yellow rectangle) whose position is unknown, reaching different targets (shown in yellow points behind the gate). The safety constraint $\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}$ is parameterized using a linear combination of spacial radial basis functions (RBFs), defined on y-z plane at $x=10$. The drone’s control inputs are the thrusts of four rotors and human correction is applied via keyboard, with specific keys mapping to specific correction commands to the drone’s control space. See the detailed task setup and keyboard interface in Appendix B-C.
(2) Franka arm reaching game. As in Fig. 16, the goal of the Franka arm reaching game is to teach a Franka arm such that its gripper hand successfully reaches different target positions (cyan ball) through an unknown narrow slot (two yellow bars). Note that the safety constraint is only applied on the end effector and the collision between the link and the slot is ignored. The safety constraint function $\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}$ is parameterized as a weighted combination of RBFs defined on the $z$ axis and rotations around the $x$ axis. Human corrections are applied via keyboard, with specific keys mapping to specific correction commands applied to the robot control space $\mathbb{R}^{6}$ (we use operational space control). See the detailed task setup and keyboard interface in Appendix B-D.
In both games, the robots are controlled with MPC controller (3) but fail to maintain safety with the initial safety constraints. To simulate an emergency stop, the “Enter” key is used as a reset command, allowing users to reset the environment to its initial state whenever the participant deems the robot violating the safety constraint.
(a) Tracking View
(b) Fixed View
Figure 15: Drone navigation game. (a) Tracking view. (b) Fixed view. Users can see the motion of the drone with the help of these two views. The $xyz$-axis (red, green, and blue axes), the gate (yellow bars), the drone, the local plan solved from the current MPC (red curve) and the motion target (yellow balls) are visible to users. In each game reset, the user will see one target randomly.
(a) Starting Configuration
(b) Motion Illustration
Figure 16: Franka arm reaching game. (a) The starting pose of the arm. RGB arrows denote the world $xyz$-frame. (b) One example of successful reaching, and long green arrow denotes the motion direction. Users can see the motion of the robot arm by adjusting the view. The slot, the robot arm, and current motion target (cyan ball) are all visible to users.
(a) After 2 corrections
(b) After 6 corrections
(c) After 12 corrections
Figure 17: Illustrations of the drone MPC rollout trajectory after $2,6,12$ corrections. The red stars denote the targets and blue curves denote the trajectories. The gate is drawn as black frame and the collision is highlighted in green crosses. User corrections typically lower the drone to make it pass through the gate.
VII-B Participants, Procedures, and Metrics
VII-B1 Participants
We recruited 15 participants from Arizona State University, including undergraduates, masters, and PhD students with different backgrounds. Among all participants, 13 are male and 2 female. All participants are new to the user study. This user study had been reviewed and approved by the Institutional Review Board (IRB) of Arizona State University, and all participants had signed the consent forms.
VII-B2 Procedure
Each participant was instructed to play the previously developed two simulation games. In each game, each participant plays 10 trials.
In the drone navigation game, a trial is deemed successful when the drone navigates through the gate without collision and reaches different targets 6 consecutive times with the learned safe MPC policy. A trial is deemed failed when there are too many corrections (exceeding 25), or the user believes the robot cannot be corrected anymore and gives up.
In the Franka reaching game, a trial is deemed successful when the robot arm reaches different targets with the end-effector having no collision with the slot. The failure of one trial is determined when there are too many corrections (exceeding 25) or the user believes the robot cannot be corrected anymore and gives up.
The result of every trial is recorded. Before formally starting the study, every user is instructed about the settings of the game and is given 2 chances to interact with the environment hands-on to get familiar with the keys.
VII-B3 Metrics for Evaluation
Two major metrics are used to evaluate the performance of the proposed algorithm:
  • (I)
    Success rate: The overall and individual success rate suggest the effectiveness of the proposed method.
  • (II)
    Numbers of corrections needed for a successful trial: The averaged number of corrections performed by users in successful trials is a measure for the efficiency of the proposed method. Fewer corrections in successful trials suggest better human data efficiency.
VII-C Result and Analysis
VII-C1 User Study Metrics Report
TABLE II: Result of user study for simulation games
Environment Success Rate Correction count
Drone navigation $89.3\%\pm 7.7\%$ $15.7\pm 10.6$
Franka arm reaching $89.3\%\pm 11.2\%$ $16.5\pm 6.1$
For the user study, we report the success rate and correction count in a successful game trial in Table II. Both metrics are averaged across users. The success rate is $89.3\%\pm 7.7\%$ for the drone navigation game and $89.3\%\pm 11.2\%$ for the Franka arm reaching game. The number of corrections for a successful trial is $15.7\pm 10.6$ for the drone navigation game and $16.5\pm 6.1$ Franka arm reaching game. In both games, each learning update of $\boldsymbol{\theta}$ takes approximately 0.15 seconds, during which the simulation briefly pauses. This brief pause is imperceptible to the user, especially since human corrections occur infrequently.
The above result illustrates that the proposed method is effective (i.e., with a high success rate) in learning a safety constraint for varying tasks (recall targets are changing across trials). It also shows the efficiency of the proposed method — successful learning only requires tens of human corrections. Since this method is online and the user can perform multiple corrections in one robot rollout trajectory, the constraint can be learned in a few runs. In Table II, one can note that the drone navigation task has a higher variance in the correction count than the Franka arm reaching task. This could be because the drone game has a larger motion range and longer task rollout horizon. As a result, human correction strategies, such as correction position and timing, can vary significantly. This correction diversity leads to a larger variance of correction count.
VII-C2 Analysis for Learned Constraint and Robot Motion
To illustrate learning progress, in Fig. 17 and Fig. 18, we show the intermediate robot motion at different numbers of human corrections. Fig. 17 shows the trajectory of the drone after different numbers of human corrections. Fig. 18 shows the pose of the robot gripper when passing the slot after a different number of human corrections.
(a) After 4 Corrections
(b) After 17 Corrections
Figure 18: Screenshots of the robot gripper when passing through the slot after $k=4,17$ corrections, given three different targets (cyan ball). (a): After 4 corrections, the gripper collides with the slot. (b): After 17 corrections, the gripper succeeds in passing the slot for all targets.
To better understand the learned safety constraint ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})$, in Fig. 19, we draw the heat map of the learned ${g}_{\boldsymbol{\theta}^{*}}(\boldsymbol{\xi})$ for the drone navigation game, with $\boldsymbol{\theta}^{*}$ the averaged of weights learned. Recall in drone navigation game, ${g}_{\boldsymbol{\theta}}$ is defined on y-z plane at $x=10$ (true gate position is $x_{\text{gate}}=9.8$). In the heat map, we draw 00-level contour of ${g}_{\boldsymbol{\theta}^{*}}$ in dashed lines, which is the learned safety boundary. Also, we plot the $yz$-plane projection of the gate, which is the black box.
Fig. 19 shows that the learned safety boundary (corresponding to ${g}_{\boldsymbol{\theta}^{*}}(\boldsymbol{\xi})\leq 0$) is only similar to the top edge of the gate. This is explained by the following reasons. As shown in Fig. 15, the targets are set higher than the gate, and thus during the learning process, the motion of the drone is always going up, as shown in Fig. 17. As a result, the majority of human corrections happen when the drone approaches the top edge of the gate, leaving the learned safety boundary accurate near the top edge of the gate and bottom edge untouched. It is expected that to learn a safety constraint that fully represents the gate, lower target positions should be added so that the drone will receive human corrections near the bottom edge of the gate.
Figure 19: Heatmap for the learned constraint in the drone game. The black solid line denotes the actual position of the gate in $yz$-plane. The black dashed curve denotes the boundary of the learned safety zone. Since the targets are higher than the gate (see Fig. 15), during the learning process, the motion of the drone is always going up, as shown in Fig. 17. As a result, the majority of human directional corrections happen when the drone approaches the top edge of the gate, leaving the learned safety boundary accurate near the top edge of the gate and bottom boundary untouched.
VIII User Study in Real-world Experiment
In this section, we conduct a user study to evaluate the performance of our safe MPC alignment algorithm in a hardware Franka Research 3 experiment. We focus on the task of mobile liquid pouring, i.e., the robot performs liquid pouring while moving from one location to targets. In certain scenarios, this kind of task can be safety-critical: for example, if the liquid is corrosive/chemical/toxic, inappropriate mobile pouring will cause safety concerns. The task is challenging because it is hard to manually define a safety constraint for the robot to accomplish liquid pouring without spilling. The Franka robot arm will learn a safe MPC controller from human physical corrections.
VIII-A Hardware Setup
[…]
The setup of mobile water pouring task is shown in Fig. 20. A water pot is gripped by the robot end-effector. The robot moves from the starting point $[0.22,-0.5,0.48]^{\mkern-1.5mu\mathsf{T}}$ (with quaternion $[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$) to target point $[-0.65,-0.5,0.5]^{\mkern-1.5mu\mathsf{T}}$ (with quaternion $[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$). During the motion, the robot needs to pour more than 50 grams of water into a bar sink. The bar sink is placed in the middle of the robot path, but the precise size and location of the bar sink are unknown.
Figure 21: System diagram for the hardware implementation. Different blocks indicate separate processes connected via arrows that indicate communication via ROS.
As shown Fig. 21, the hardware Franka arm platform has three main components: the parameterized safe MPC controller (high-level controller), the low-level controller, and safe MPC alignment module. The safe MPC controller follows the safe MPC formulation of Franka arm reaching game in Appendix B-D (recall that the learnable safety constraint is parameterized by RBFs). The parameters of this Safe MPC controller are listed in Table III. The safe MPC controller receives the end-effector’s current pose $\boldsymbol{x}$ and generate the action $\boldsymbol{u}$ (see (58) in Appendix), which is the desired velocity of the end-effector. The low-level controller implements the operational space control to track the $\boldsymbol{u}$ given by the safe MPC controller.
The safe MPC alignment module updates the parameterize $\boldsymbol{\theta}$ of the safe MPC controller. The update only happens when the alignment module detects a human correction $\boldsymbol{a}$. The module implements our safe MPC alignment algorithm. Note that the alignment module and controller modules runs different process, thus the learning update does not affect the real-time MPC control of the robot arm. ROS is used for communication between different processes.
TABLE III: parameter list in real-world user study
Parameter Value
$Q_{r}$ $\mathrm{diag}(0.8,0.8,0.8)$
$k_{q}$ 0
$R_{v}$ $\mathrm{diag}(30,30,1)$
$R_{w}$ $\mathrm{diag}(0.85,0.85,0.85)$
$S_{r}$ $\mathrm{diag}(80,80,80)$
$\mu_{q}$ 80
In the safe MPC controller, the cost of the MPC controller is set to penalize the current and target, as in Section B-D. Since it is difficult to explicitly define a parameterized safety constraint using the state of pouring water, we instead define the constraints only using the robot state. Specifically, we use the weighted RBFs to parameterize the safety constraints constraint, similar to Appendix B-D.
As shown in Fig. 20, to give a physical correction to the Franka robot arm, a human user need physically touch/hold any link of the robot arm, and the applied force will immediately activate the zero-torque control mode. Suppose the joint positions of the robot arm before and after human physical correction are $\mathbf{q}_{\text{before}}$ and ${\mathbf{q}_{\text{after}}}$, respectively, we calculate the corresponding raw human correction $\boldsymbol{a}_{r}$ using
$\boldsymbol{a}_{r}=\mathrm{J}(\mathbf{q}_{\text{before}})({\mathbf{q}}_{\text{after}}-\mathbf{q}_{\text{before}}),$ (35)
with $\mathrm{J}$ the Jacobian of the end-effector. The calculated $\boldsymbol{a}_{r}\in\mathbb{R}^{6}$ is considered as the displacement of the end-effector pose, intended by the human user.
(a) After 2 Corrections
(b) After 3 Corrections
(c) After 5 Corrections
Figure 22: Illustrations of one trial of the real-world experiment. The trajectories of the watering pot is denoted by blue dashed lines. (a) After 2 corrections, the arm begins tilting the pot but spills out the water. (b) After 3 corrections, the arm can pour water into the sink but the human is not satisfied with its height and altitude of pouring. (c) After 5 corrections, the arm learns to tilt only along the direction of the $x-$axis and successfully aligns with human intent.
VIII-B Participants, Procedures, and Metrics
All participants recruited for the previous two simulation games in Section VII-B also participate the hardware experiment. They are told of the objective before starting trials. One episode is defined as one robot motion from the starting position to the targets, during which the human can apply corrections at any time instances. One learning trial may include multiple episodes (with a maximum of 15 corrections). During learning, the nozzle of the water pot is sealed, and we switch to the unsealed nozzle and measure the amount of water poured into the flower pot after the user is satisfied with the motion of the robot motion. The success learning trial is defined when the amount of poured water is over 50g. Each user performs 5 trials, and the same metrics in Section VII-B are recorded.
VIII-C Result and Analysis
In the real-world robot experiment, the success rate is $84.0\%\pm 16.7\%$ and the average number of corrections needed for successful trials is $6.88\pm 3.26$. The results show the efficacy and efficiency of our method in real-world robot learning applications. To illustrate the robot learning progress, in Fig. 22, we use one trial to show the robot behavior at different numbers of human corrections.
Since this real-world Franka arm experiment is similar to the simulated Franka arm reaching game in Appendix B-D, we compare the user study results for both and have the following conclusion and analysis.
(I) The real-world experiment has a slightly lower success rate: While the difference between simulations and real-world experiments is small in success rate, we have identified some factors during our experiment that may contribute to such lower success rate, including both the human aspects and hardware aspects. On human aspects, most of the participants have no experience or knowledge in operating robot hardware, and thus when directly interacting with hardware, their corrections can be conservative or ineffective due to safety considerations. On hardware side, noisy correction signals and the possible timing delay between the recorded human correction and robot state can contribute to the failure of the learning.
(II) The real-world experiment has a less number of human corrections: This could be attributed to two reasons. First, the constraint needed for accomplishing the water pouring task could be more flexible than that for reaching task in the simulation game Appendix B-D. This makes the safety constraints easy to learn for the real-world task. Second, different from simulation games, where human users can only perceive the robot arm from the screen, a human user has a better perception of the robot’s motion in real world. A good perception of the robot’s state could help humans give more effective feedback each time, reducing the total number of human interactions for successful robot learning.
VIII-D Discussion
Based on the observations in both the hardware and simulation user studies, we discuss some design choices and potential limitations of the proposed method.
VIII-D1 Learning Task-Independent Safety Constraints
In the real-world and simulation experiments, due to the insufficiency of task sampling, the learned safety constraint could be task-specific, instead of task-independent. For instance, in the drone navigation game, since the three task targets are placed above the gate, the learned constraint shown in Fig. 19 can only reflect the upper edge of the unknown gate. One limitation of this is that the learned task constraints have poor generalization to new tasks. For example, in drone navigation, if a new target is placed below the gate, the drone with the learned safety constraint Fig. 19 will fail.
To improve the generalization of the learned safety constraints or learning task-independent safety constraints, one solution is to increase the diversity of task samples. By doing so, it is expected that the robot can receive human correction on the entire boundary of the safety region, making the learned safety constraints reusable in new tasks.
VIII-D2 Robustness Against Human Mistakes
The proposed safe MPC alignment is based on the hypothesis cutting. In our formulation, a hypothesis cut (7) is based on an average of multiple human corrections, such an averaging treatment can mitigate human mistakes for one cut. But still, in practice, human irrationality can lead to the wrong cut of the hypothesis space, which is the main reason for failure in user studies. For example, in the drone navigation game, we found that failure game trials usually happen when a human gives too many wrong directional corrections. A failure case is in Fig. 23.
Figure 23: One illustration of the failure case in the drone game user study. A human user makes a false "up" correction at the beginning and fails to correct the motion of the drone.
To further improve the robustness of the method against human mistakes, two extensions of the current method can be explored. The first is to add a relaxation factor $\epsilon\geq 0$ in our cutting hypothesis. For example, we can re-write (7) as
$\forall\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H},\ \operatorname{\mathbb{E}}_{\mathbf{a}}\left\langle\mathbf{a},-\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})\right\rangle+\epsilon\geq 0.$ (36)
Another is to add a selection mechanism on human correction. For example, one can use some statistical method to rule out mistake corrections before performing a hypothesis cut. We will leave those possible extensions in our future work.
IX Conclusion
This paper proposes safe MPC alignment method to online learn a safety constraint in safe MPC controllers from human directional feedback. The method, named hypothesis space cutting, has geometric interpretation: it exponentially cuts the hypothesis space based on each received human directional feedback. The method has a provable guarantee for the total number of human feedback needed for learning, and also can declare hypothesis misspecification. Numeric examples, user studies on simulation games, and a user study on the real-world robot arm, covering varying robot tasks, were used to validate the effectiveness of the proposed approach. The results show that the proposed method can learn safety constraints effectively and efficiently. To our knowledge, this is the first certifiable method to enable online learning of safety constraints with human feedback.
Appendix A Proof of lemmas and theorems
A-A Proof of Lemma 2
Proof.
We prove this lemma by induction. First, we have $\bar{\Theta}_{H}\subset\Theta_{0}$. Suppose $\bar{\Theta}_{H}\subset\Theta_{i-1}$, $i=1,2,3,...$. Due to the condition $\bar{\Theta}_{H}\subset\mathcal{C}_{i}$ in (12), following the update (13), we have $\bar{\Theta}_{H}\subset\Theta_{i}=\Theta_{i-1}\cap\mathcal{C}_{i}$, which completes the proof. ∎
A-B Proof of Lemma 3
Proof.
(14) can be directly obtained from (12), (2) and (5). Based on the definition of the penalty objective function (2), $\mathbf{\bar{a}}_{i},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta})\right>\leq 0$ is a linear fractional inequality
$\mathbf{\bar{a}}_{i},\nabla J(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>+\gamma\frac{1}{-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})}\left<\mathbf{\bar{a}}_{i},\nabla g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>\leq 0,$ (37)
with a positive denominator $-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})>0$ due to (12).
Multiplying the denominator on both side and substituting the weight-feature expression of $g_{\boldsymbol{\theta}}$ in (5) yields $\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}\leq b_{i}$ in Lemma 3. The condition $-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})>0$ itself yields $\ {\boldsymbol{\theta}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<-\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})$ in Lemma 3.
This completes the proof. ∎
A-C Proof of Lemma 4
Proof.
For any $i=1,2,3,...$, the plan $\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ is a solution to the optimization (3), with $\boldsymbol{\theta}=\boldsymbol{\theta}_{i}$. Recall Lemma 1, $g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<0$. Extending the feature-weight expression of $g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})$, one can get $\boldsymbol{\theta}_{i}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<-\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})$. Since the optimal solution $\boldsymbol{u}^{\boldsymbol{\theta}_{i},\gamma}_{0:T-1}$ is a minimizer to (3), it means
$\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta}_{i})=0.$ (38)
Note that with $g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<0$, the above $\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta}_{i})$ is well defined. With (38),
$\mathbf{a},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta}_{i})\right>=0,$ (39)
always holds. Extending the above (39) leads to
$\mathbf{a},\nabla J(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>-\gamma\frac{1}{g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})}\left<\mathbf{a},\nabla g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>=0.$ (40)
Due to $-g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})>0$, one can multiple $-g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})$ on both sides of the above equality, which leads to $\boldsymbol{\theta}_{i}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}=b_{i}$ based on the definition (15). The second term in (18) directly follows $g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<0$. This concludes the proof. ∎
A-D Proof of Lemma 5
Proof.
By Lemma 4, the MVE center $\boldsymbol{\theta}_{i}=\bar{\boldsymbol{d}}_{i-1}$ is on the cutting hyperplane $\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}=b_{i}$. Directly from [63], one has
$\frac{\mathbf{Vol}(\Theta_{i-1}\cap\{\boldsymbol{\theta}|\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}\leq b_{i}\})}{\mathbf{Vol}(\Theta_{i-1})}\leq 1-\frac{1}{r}.$ (41)
Since
$\mathbf{Vol}(\Theta_{i-1}\cap\{\boldsymbol{\theta}|\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}\leq b_{i}\})\geq\mathbf{Vol}(\Theta_{i-1}\cap\mathcal{C}_{i})=\mathbf{Vol}(\Theta_{i}),$ (42)
based on (12), (13), one can get
$\frac{\mathbf{Vol}(\Theta_{i})}{\mathbf{Vol}(\Theta_{i-1})}\leq 1-\frac{1}{r},$ (43)
which concludes the proof. ∎
A-E Proof of Theorem 1
Proof.
We complete the proof in two parts. First, we prove the number of learning iterations is less than $K$, then we use this to prove there will be a $\boldsymbol{\theta}_{k}\in\bar{\Theta}_{H}$, $k=0,1,2...,K-1.$
We first prove the total number of iterations must be less than $K$ by contradiction. Suppose the learning iteration number $i\geq K$. With recursively applying lemma 5, one can obtain
$\mathbf{Vol}(\Theta_{i})\leq(1-\frac{1}{r})^{i}\mathbf{Vol}(\Theta_{0}).$ (44)
When $i\geq K$,
$\mathbf{Vol}(\Theta_{i})<\frac{\tau_{r}\rho_{H}^{r}}{\mathbf{Vol}(\Theta_{0})}\mathbf{Vol}(\Theta_{0})=\tau_{r}\rho_{H}^{r}\leq\mathbf{Vol}(\bar{\Theta}_{H})$ (45)
However, from Lemma 2, one has $\bar{\Theta}_{H}\subset\Theta_{i}$, meaning $\mathbf{Vol}(\bar{\Theta}_{H})\leq\mathbf{Vol}(\Theta_{i})$. This leads to a contradiction with (45). Therefore, the total number of iterations must be less than $K$.
Next, we prove $\exists k<K$ such that $\boldsymbol{\theta}_{k}\in\bar{\Theta}_{H}$. In contradiction, suppose $\forall k<K$, $\boldsymbol{\theta}_{k}\notin\bar{\Theta}_{H}$. By Lemma 2, this means that for all $i<K$, $\bar{\Theta}_{H}\subset\mathcal{C}_{i}$. Then, the algorithm will continue for $i=K$. Consequently, the total number of iterations can reach $K$, which contradicts that the algorithm will terminate within $K$ iterations. This concludes the proof. ∎
A-F Proof of Theorem 2
Proof.
We complete the proof in two parts. First we prove the algorithm converges to $\Theta^{*}$ which has $\mathbf{Vol}(\Theta^{*})=0$. Based on this, we next prove $\Theta^{*}\cap\mathbf{int}\ \Theta_{0}=\emptyset$.
We first prove the algorithm converges to $\Theta^{*}$ which has $\mathbf{Vol}(\Theta^{*})=0$. Given the condition $\Theta_{0}\cap\bar{\Theta}_{H}=\emptyset$, the algorithm will never terminate because $\forall i=1,2,3,...\infty$, $\mathcal{C}_{i}\supset\bar{\Theta}_{H}$ due to Lemma 2. Since the volume of the initial hypothesis space is finite, by applying Lemma 5 we have
$\mathbf{Vol}(\Theta^{*})=\lim_{i\rightarrow\infty}(1-\frac{1}{r})^{i}\mathbf{Vol}(\Theta_{0})=0.$ (46)
Next, we prove $\Theta^{*}\cap\mathbf{int}\ \Theta_{0}=\emptyset$ by contradiction. Suppose there exists $\widehat{\boldsymbol{\theta}}\in\Theta^{*}\cap\mathbf{int}\ \Theta_{0}$. Define the ball contained in $\bar{\Theta}_{H}$ to be $\mathcal{B}({\boldsymbol{\theta}^{c}_{H}},{\rho_{H}})=\{\boldsymbol{\theta}\ |\ \left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}^{c}\right\rVert\leq\rho_{H}\}\subseteq\bar{\Theta}_{H}$ with $\boldsymbol{\theta}_{H}^{c}$ as the center. By the equation (13), we can note
$\Theta^{*}=\Theta_{0}\cap{\mathcal{C}},$ (47)
with ${\mathcal{C}}=\bigcap_{i=1}^{\infty}\mathcal{C}_{i}$ a convex set. So we have
$\Theta^{*}\subseteq\mathcal{C}.$ (48)
By equation (12),
$\mathcal{B}_{H}\subseteq\bar{\Theta}_{H}\subseteq\mathcal{C}.$ (49)
From the relationship between sets above, $\boldsymbol{\theta}_{H}^{c}\in\mathcal{C},\widehat{\boldsymbol{\theta}}\in\mathcal{C}$.
Since $\boldsymbol{\theta}_{H}^{c}\notin\Theta_{0}$, there exists $\mu>0$ such that
$\widetilde{\boldsymbol{\theta}}=\widehat{\boldsymbol{\theta}}+\mu(\boldsymbol{\theta}_{H}^{c}-\widehat{\boldsymbol{\theta}})\in\mathbf{int}\ \Theta_{0}.$ (50)
Then, one can define a ball $\mathcal{B}(\widetilde{\boldsymbol{\theta}},d)$ centered at $\widetilde{\boldsymbol{\theta}}$ with radius
$d=\min_{\boldsymbol{\theta}\in\mathbf{bd}\ \Theta_{0}}\left\lVert\widetilde{\boldsymbol{\theta}}-\boldsymbol{\theta}\right\rVert.$ (51)
Thus, $\mathcal{B}(\widetilde{\boldsymbol{\theta}},d)\subset\mathbf{int}\ \Theta_{0}$.
Next, we will show $\mathcal{B}(\widetilde{\boldsymbol{\theta}},\mu\rho_{H})\subset\mathcal{C}$. In fact, for any $\boldsymbol{\theta}$ satisfying $\left\lVert\boldsymbol{\theta}-\widetilde{\boldsymbol{\theta}}\right\rVert\leq\mu\rho_{H}$, one has
$\begin{split}\boldsymbol{\theta}&=\mu(\frac{1}{\mu}(\boldsymbol{\theta}-\widetilde{\boldsymbol{\theta}})+\boldsymbol{\theta}_{H}^{c})+\widetilde{\boldsymbol{\theta}}-\mu\boldsymbol{\theta}_{H}^{c}\\ &=\mu\underbrace{(\frac{1}{\mu}(\boldsymbol{\theta}-\widetilde{\boldsymbol{\theta}})+\boldsymbol{\theta}_{H}^{c})}_{\in\mathcal{B}_{H}\subseteq\mathcal{C}}+(1-\mu)\underbrace{\widehat{\boldsymbol{\theta}}}_{\in\Theta^{*}\subseteq\mathcal{C}}\end{split}.$ (52)
The convexity of $\mathcal{C}$ leads to $\boldsymbol{\theta}\in\mathcal{C}$, which means the ball $\mathcal{B}_{\mu}\triangleq\{\boldsymbol{\theta}\ |\ \left\lVert\widetilde{\boldsymbol{\theta}}-\boldsymbol{\theta}\right\rVert\leq\mu\rho_{H}\}\subset\mathcal{C}$.
Let a new ball be $\mathcal{B}(\widetilde{\boldsymbol{\theta}},\min{(\mu\rho_{H},d)})$. Based on the previous two results of $\mathcal{B}(\widetilde{\boldsymbol{\theta}},d)\subset\mathbf{int}\Theta_{0}$ and $\mathcal{B}(\widetilde{\boldsymbol{\theta}},\mu\rho_{H})\subset\mathcal{C}$, one has $\mathcal{B}(\widetilde{\boldsymbol{\theta}},\min{(\mu\rho_{H},d)})\subset(\Theta_{0}\cap\widetilde{\mathcal{C}})=\Theta^{*}$. As a result,
$\mathbf{Vol}(\Theta^{*})\geq\mathbf{Vol}\,\,\mathcal{B}(\widetilde{\boldsymbol{\theta}},\min{(\mu\rho_{H},d)})>0$ (53)
which contradicts $\mathbf{Vol}(\Theta^{*})=0$ from (46). Thus, $\Theta^{*}\cap\mathbf{int}\ \Theta_{0}=\emptyset$ must hold. This completes the proof. ∎
Appendix B Settings of Experiments and User Studies
B-A Inverted Pendulum
In the dynamics of an inverted pendulum (26), the physical parameters is set in Table IV.
TABLE IV: Physics Parameters in Pendulum Numerical Study
Parameter Value Interpretation
$m$ 1kg Weight of the pendulum
$l$ 1m Length of the pendulum
$\mathrm{g}$ $10\mathrm{m/s}^{2}$ Gravitational Acceleration
$d$ 0.1 Dampening Coefficient
When the system state violates ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})\leq 0$ or it is near the target, the pendulum state is uniformly randomly reset between $[0,0]^{\mkern-1.5mu\mathsf{T}}$ and $[\frac{2\pi}{3},3]^{\mkern-1.5mu\mathsf{T}}$. Simulated human correction will continue to be applied until the learned parameters $\boldsymbol{\theta}_{i}$ fall in $\bar{\Theta}_{H}$, which terminates one trial.
B-B Two-link Robot Arm
We assume each links has unit length and mass, and the mass is evenly distributed along the link. The corresponding inertia matrix and Coriolis term can then be calculated.
B-C Drone Navigation Game
B-C1 Environment
The drone starts from $[0.1,y_{\text{init}},0.5]^{\mkern-1.5mu\mathsf{T}}$ with $y_{\text{init}}$ uniformly randomly in an interval $[4,6]$. The starting quaternion $\boldsymbol{q}_{\text{init}}=[1,0,0,0]^{\mkern-1.5mu\mathsf{T}}$. We set three different targets at positions $[19,1,9]^{\mkern-1.5mu\mathsf{T}}$, $[19,5,9]^{\mkern-1.5mu\mathsf{T}}$ and $[19,9,9]^{\mkern-1.5mu\mathsf{T}}$, with quaternion $[1,0,0,0]^{\mkern-1.5mu\mathsf{T}}$. For the narrow gate (yellow rectangle) in Fig. 15, both its size and position are unknown, and one only roughly knows its x position $x_{\text{gate}}\approx 9.8$.
Since the location of the gate is unknown, the drone needs to learn a safety constraint in its safe MPC policy with humans guidance in order to successfully fly through the gate. In the drone MPC policy (1), the dynamics and its parameters are set following ([20]); and we generically set the constraint function ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})$ using a set of $N$ Gaussian radial basis functions (RBFs)
${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})=\phi_{0}+\sum\nolimits_{i=1}^{N}\theta_{i}\phi_{i}(\boldsymbol{\xi}).$ (54)
As discussed in Section III-C4, we fix $\phi_{0}=-1$ to avoid learning ambiguity, and use $N=20$ temporally accumulated RBFs defined as
$\phi_{i}(\boldsymbol{\xi})=\sum\nolimits_{t=0}^{T}\beta_{t}\mathrm{RBF}_{i}(\boldsymbol{x}_{t}),$ (55)
where $\beta_{t}$ is a weighting factor to balance the immediate and future constraint satisfaction. If one emphasizes instant safety, more weights will be given to the initial time step. In our implementation, we choose an exponentially decaying $\beta_{t}$ starting at $t=5$, i.e. $\beta_{t}=0,\ t<5$ and $\beta_{t}=0.9^{t-5},\ t\geq 5$ to achieve better predictive safety performance. In (55), each RBF is defined as
$\mathrm{RBF}_{i}(\boldsymbol{x}_{t})=\begin{cases}\exp(-\epsilon^{2}\left\lVert r_{xy,t}-c_{xy,i}\right\rVert^{2}),\,\,i=1,...,10\\ \exp(-\epsilon^{2}\left\lVert r_{xz,t}-c_{xz,i}\right\rVert^{2}),\,\,i=11,...,20\end{cases}.$ (56)
Here, the variables $r_{xy,t}=[r_{x,t},r_{y,t}]^{\mkern-1.5mu\mathsf{T}}$ and $r_{xz,t}=[r_{x,t},r_{z,t}]^{\mkern-1.5mu\mathsf{T}}$ with $r_{x},r_{y},r_{z}$ from $\boldsymbol{x}_{t}$. The $i$th center $c_{xy,i}=[c_{x,i},c_{y,i}]^{\mkern-1.5mu\mathsf{T}}$ or $c_{xz,i}=[c_{x,i},c_{z,i}]^{\mkern-1.5mu\mathsf{T}}$ is defined with $c_{x,i}={x_{\text{gate}}}=9.8$, $c_{y,i}$ and $c_{z,i}$ are evenly chosen on interval $[0,10]$. We set $\epsilon=0.45$. The initial hypothesis space $\Theta_{0}$ for possible $\boldsymbol{\theta}$ in (54) is set as $\boldsymbol{c}_{l}=[-80,\dots,-80]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{u}=[100,\dots,100]^{\mkern-1.5mu\mathsf{T}}$. $\gamma$ is 50. This initialization blocks the drone from advancing to the gate at the beginning of the game, leading to conservative behavior in the learning process. The drone will proceed toward the gate after few corrections.
B-C2 Human Correction Interface
The motion of the drone is rendered in two camera views in MuJoCo visualization (Fig. 15) for better human perception: one is the first-person view tracking the drone, the other a third-person fixed view. The gate, the current MPC planning solution, and the target position are visible to the users.
Human users apply corrections to the drone via a keyboard during the drone flying. The user is informed of the mappings between the key press and the corresponding correction in the game, listed in Table V. The safety constraint (54) in the drone’s MPC policy is immediately updated once a user correction is detected. When the user presses "ENTER", the drone is reset to the initial position and target will randmize.
TABLE V: keyboard commands in uav user study
Key Correction Interpretation
"up" $[-1,0,1,0]^{\mkern-1.5mu\mathsf{T}}$ Go forward (adding positive y-axis torque)
"down" $[1,0,-1,0]^{\mkern-1.5mu\mathsf{T}}$ Go backward (adding negative y-axis torque)
"left" $[0,1,0,-1]^{\mkern-1.5mu\mathsf{T}}$ Go left (adding negative x-axis torque)
"right" $[0,-1,0,1]^{\mkern-1.5mu\mathsf{T}}$ Go right (adding positive x-axis torque)
"shift" $[1,1,1,1]^{\mkern-1.5mu\mathsf{T}}$ Go up (increasing all thrusts)
"ctrl" $-[1,1,1,1]^{\mkern-1.5mu\mathsf{T}}$ Go down (decreasing all thrusts)
"enter" N/A Reset robot position and switch target.
B-D Franka Constrained Reaching Game
B-D1 Environment
The arm starts at a configuration in which the end-effector position $\boldsymbol{r}_{\text{init}}=[0.22,-0.5,0.48]^{\mkern-1.5mu\mathsf{T}}$ and altitude $\boldsymbol{q}_{\text{init}}=[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$. The reaching target position $\boldsymbol{r}^{*}$ will be randomly chosen from three options $[-0.6,-0.5,0.4]^{\mkern-1.5mu\mathsf{T}}$, $[-0.6,-0.5,0.5]^{\mkern-1.5mu\mathsf{T}}$, $[-0.6,-0.5,0.6]^{\mkern-1.5mu\mathsf{T}}$. The reaching target attitude is $\boldsymbol{q}^{*}=[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$. The slot size and position are unknown except its rough x-position which is $x_{\text{slot}}\approx-0.15$.
We adopt a two-level control scheme, where a safe MPC policy (high-level) is defined for the end effector, which generates the desired velocity for a given low-level operational space controller to track. In the safe MPC policy, we model the end-effector dynamics as
$\begin{gathered}\dot{\boldsymbol{r}}=\boldsymbol{v}^{d},\\ \dot{\boldsymbol{q}}=\frac{1}{2}\Omega(R^{\mkern-1.5mu\mathsf{T}}(\boldsymbol{q})\boldsymbol{w}^{d})\boldsymbol{q},\end{gathered}$ (57)
where $\boldsymbol{r}$ and $\boldsymbol{q}$ is the position and the quaternion of the arm end-effector in the world frame. $R^{\mkern-1.5mu\mathsf{T}}(\boldsymbol{q})$ is the rotation matrix from the end-effector frame to the body frame. $\Omega(\cdot)$ is the matrix operator of the angular velocity for quaternion multiplication. Define the robot state and control input
$\boldsymbol{x}_{t}=[\boldsymbol{r}_{t},\ \boldsymbol{q}_{t}]\in\mathbb{R}^{7},\quad\boldsymbol{u}_{t}=[\boldsymbol{v}^{d}_{t},\ \boldsymbol{w}^{d}_{t}]\in\mathbb{R}^{6},$ (58)
respectively.
The above end-effector dynamics is discretized with $\triangle t=0.1\mathrm{s}$. In (1), the MPC horizon is $T=20$. Let $\boldsymbol{x}^{*}=[\boldsymbol{r}^{*}\ \boldsymbol{q}^{*}]$ denote the target end-effector pose. The step cost $c(\boldsymbol{x}_{t},\boldsymbol{u}_{t})$ and terminal cost $h(\boldsymbol{x}_{T})$ in (1) is set to
$c(\boldsymbol{x}_{t},\boldsymbol{u}_{t})=$ $(\boldsymbol{r}_{t}-\boldsymbol{r}^{*})^{\mkern-1.5mu\mathsf{T}}Q_{r}(\boldsymbol{r}_{t}-\boldsymbol{r}^{*})+k_{q}\mathbf{diff}(\boldsymbol{q}_{t},\boldsymbol{q}^{*})$ (59)
$\quad\quad+\boldsymbol{v}_{I,t}^{\mkern-1.5mu\mathsf{T}}R_{v}\boldsymbol{v}_{I,t}+\boldsymbol{w}_{I,t}^{\mkern-1.5mu\mathsf{T}}R_{w}\boldsymbol{w}_{I,t},$
$h(\boldsymbol{x}_{T})=$ $(\boldsymbol{r}_{T}-\boldsymbol{r}^{*})^{\mkern-1.5mu\mathsf{T}}S_{r}(\boldsymbol{r}_{T}-\boldsymbol{r}^{*})+\mu_{q}\mathbf{diff}(\boldsymbol{q}_{T},\boldsymbol{q}^{*}).$
In this game, the cost parameters are set as $Q=\mathrm{diag}(2,2,2)$, $k_{q}=1$, $R_{v}=\mathrm{diag}(30,30,1)$, $R_{w}=\mathrm{diag}(0.85,0.85,0.85)$, $S_{r}=\mathrm{diag}(80,80,80)$, and $\mu_{q}=60$. The output of the MPC controller is sent to the low-level controller to track.
Similar to Appendix B-C, to constrain $r_{z}$ and $\boldsymbol{q}$ for proper height and tilt in order to pass through the slot. We use a set of $N=20$ Gaussian radial basis function (RBFs) to parameter the safety constraint:
${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})=\phi_{0}+\sum\nolimits_{i=1}^{N}\theta_{i}\phi_{i}(\boldsymbol{\xi}),$ (60)
with
$\phi_{i}(\boldsymbol{\xi})=\sum\nolimits_{t=0}^{T}\beta_{t}\mathrm{RBF}_{i}(\boldsymbol{x}_{t}),\quad\text{and}$ (61)
$\mathrm{RBF}_{i}(\boldsymbol{x}_{t}){=}\begin{cases}\sigma(r_{x,t})k_{z}\exp{(-\epsilon_{z}^{2}(r_{z,t}-c_{z,i})^{2})},\,\,i{=}1,...,10\\ \sigma(r_{x,t})\exp{(-\epsilon_{q}^{2}\mathbf{diff}(\boldsymbol{q}_{t},\boldsymbol{q}_{c,i}))},\,\,i{=}11,...,20\end{cases}.$ (62)
Here, $k_{z}$ is used to balance the scaling between position and orientation $\mathrm{RBF}$s, $\sigma(r_{x,t})$ is defined as
$\sigma(r_{x,t})=\mathrm{sigmoid}(50\big{(}(r_{x,t}-x_{\text{slot}})^{2}-(\frac{w}{2})^{2}\big{)}),$ (63)
which is used to constrain the $\mathrm{RBF}_{i}$ only taking effect around $x$ range $[x_{\text{slot}}-\frac{w}{2},x_{\text{slot}}+\frac{w}{2}]$ (recall slot’s $x$ location, slot, is a approximate value). In the above definition, $w=0.3$, $k_{z}=0.05$, $\epsilon_{z}=12$ and $\epsilon_{q}=1.8$. The center of positional $\mathrm{RBF}$, ${c_{z,i}}$, is evenly chosen on interval $[0,1]$, and the center of orientation $\mathrm{RBF}$, ${\boldsymbol{q}_{c,i}}$, is generated by $(\cos{\frac{\alpha}{2}},\sin{\frac{\alpha}{2}},0,0)*\boldsymbol{q}_{\text{init}}$ with $\alpha$ evenly choosen from $[-\pi/4,3\pi/4]$ to make sure ${\boldsymbol{q}_{c,i}}$ covers all possible orientations. $\beta_{t}=0.9^{t}$ if $t<5$ and 00 otherwise. The bounds of the initial hypothesis space $\Theta_{0}$ is set to $\boldsymbol{c}_{l}=[-3,\dots,-3]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{u}=[5,\dots,5]^{\mkern-1.5mu\mathsf{T}}$. $\gamma$ is set to 1.5. This initialization leads to a conservative MPC policy, preventing the arm from approaching the target at the beginning.
B-D2 Human Correction Interface
The motion of the robot arm is rendered in the free-view mode in MuJoCo visualization (Fig. 16), and users can drag to select his/her comfortable view. The slot and the random reaching target are visible to the users.
Human users apply their corrections to the robot arm via a keyboard. Users are informed with the following key mapping between the key press and the corresponding correction in the game. The safety constraint (60) in Franka arm’s safe MPC policy is updated immediately once a user correction is detected. When user press "ENTER", the arm is reset to the initial position and the reaching target will also change.
TABLE VI: keyboard commands in robot arm user study
Key Correction Interpretation
"left" $[0,0,0,0,0,1]^{\mkern-1.5mu\mathsf{T}}$ Clock-wise rotation of the gripper
"right" $[0,0,0,0,0,-1]^{\mkern-1.5mu\mathsf{T}}$ Anti-clock-wise rotation of the gripper
"shift" $[0,0,0,-1,0,0]^{\mkern-1.5mu\mathsf{T}}$ Go up in world frame
"ctrl" $[0,0,0,1,0,0]^{\mkern-1.5mu\mathsf{T}}$ Go down in world frame
"Enter" N/A Reset robot position and switch target.